### v1

# Murphy: Multi-Turn GRPO for Self Correcting Code Generation

Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has emerged as a powerful framework for enhancing the reasoning capabilities of large language models (LLMs). However, existing approaches such as Group Relative Policy Optimization (GRPO) and its variants, while effective on reasoning benchmarks, struggle with agentic tasks that require iterative decision-making. We introduce Murphy, a multi-turn reflective optimization framework that extends GRPO by incorporating iterative self-correction during training. By leveraging both quantitative and qualitative execution feedback, Murphy enables models to progressively refine their reasoning across multiple turns. Evaluations on code generation benchmarks with model families such as Qwen and OLMo show that Murphy consistently improves performance, achieving up to a $8\%$ relative gain in pass@1 over GRPO, on similar compute budgets.
1 Introduction
“The road to wisdom? Well, it’s plain and simple to express: err and err and err again, but less and less and less.” —Piet Hein
Qwen3-1.7BQwen3-4BOLMo-2-7B-Ins02244668810102.82.82.12.13.23.26.66.62.92.97.47.4ModelsPercentage (%)($\%$)GRPOMurphy
Figure 1: Percentage change in coding problems solved by models trained with Murphy and GRPO over the base model across three models and datasets. Murphy-trained models solve up to $4.2\%$ more problems than GRPO. See Tab. 1 for details.
A growing body of work explores large language models (LLMs) as software engineering agents that interact with their environment through code execution and feedback (Shinn et al., 2023; Lingam et al., 2025; Miserendino et al., 2025). Rather than producing a single static response, these systems operate within agentic scaffolds (Yao et al., 2023; Zhou et al., 2024) that guide iterative reasoning and allow LLMs to act, observe, and improve over multiple rounds of interaction. For example, in a typical coding agnetic scaffold (Shinn et al., 2023; Lingam et al., 2025), the agent generates a solution by performing a single or series of actions, and executes it for evaluation, often through unit tests or other automated checks. When execution fails, the agent receives feedback such as error messages, stack traces, or failing test cases, and is re-prompted with the original task, its previous attempt, and the new feedback. This process continues for multiple turns until the model succeeds or reaches a fixed iteration limit. These systems highlight the growing ability of LLMs to reason and adapt through environmental feedback at inference time. In code generation tasks (Jiang et al., 2025), such feedback naturally arises from execution logs, compiler errors, or test results. However, these methods (Shinn et al., 2023; Lingam et al., 2025) remain fundamentally training-free: they improve model behavior through structured inference via feedback rather than through parameter updates.
On the other hand, training methodologies such as Reinforcement Learning with Verifiable Rewards (RLVR) have enabled a new generation of language models (OpenAI et al., 2024; DeepSeek-AI et al., 2025; Team et al., 2025; Yang et al., 2025) to exhibit strong reasoning capabilities across mathematics, coding, and general problem-solving. Recent RLVR algorithms, including Group Relative Policy Optimization (GRPO) (Shao et al., 2024) and its extensions (Yue et al., 2025; Yu et al., 2025; Xu et al., 2025a), have become dominant approaches for post-training LLMs on verifiable reasoning tasks. However, GRPO and its extensions are fundamentally designed for a single-turn setting: they optimize model behavior using an isolated prompt–response–reward tuple, with no notion of multi-turn interaction defined in their objective. Here, a turn denotes one cycle in which the model receives a prompt (which includes feedback from its previous response), reasons based on the prompt, and generates a revised output. Taken together, these limitations underscore a methodological gap: while inference-time agentic frameworks exploit iterative feedback to refine reasoning, GRPO optimizes solely from terminal rewards on a given prompt, without leveraging intermediate environmental feedback. This motivates the following key research question.
Key Research Question: How can RLVR algorithms be extended to integrate iterative feedback, and to what extent does this improve reasoning and performance within agentic frameworks?
To address this question, we propose Murphy, a novel RLVR algorithm that extends GRPO to a multi-turn setting by conditioning optimization on intermediate environmental feedback. Extending GRPO beyond a single turn is non-trivial: it requires defining how rewards obtained in later turns should be propagated backward to earlier attempts, so that intermediate reasoning and output turns that initially failed but ultimately led to success through feedback receive appropriate credit. At the first turn, Murphy generates $G_{1}$ generations per prompt and computes group-based rewards as in GRPO. In subsequent turns, generations that fail to achieve the maximum reward, e.g., those failing unit tests or producing incorrect outputs, are refined using signals from the environment such as executor logs or test results. This feedback is appended to the original prompt, and the model is re-prompted to generate a new batch of $G_{s}$ generations (where $s$ denotes the turn) conditioned on the combined context (prompt, previous output, and feedback), repeating this process for a fixed number of turns. Rewards from successful final-turn rollouts are then propagated backward to earlier turns using Murphy’s credit-assignment criterion, allowing partially correct but improving attempts to receive credit. To manage the computational cost of multi-turn updates, Murphy employs pruning strategies that retain only promising trajectories while bounding total gradient updates per turn. In summary, our main contributions are:
Main Contributions. 1. We introduce Murphy, a multi-turn RLVR algorithm that extends GRPO to leverage execution feedback for grounded reasoning and self-correction in code generation tasks. (Sec. 4) 2. We design pruning mechanisms that enable scalable multi-turn RLVR training (Murphy) while preserving performance. (Subsec. 4.1) 3. We evaluate Murphy on three code generation benchmarks across two model families (OLMo, Qwen) and sizes (1.7B–7B), achieving up to $5\%$ pass@1 over GRPO baselines. (Sec. 5)
2 Related Work
LLM Agents for Software Development.
Recent studies (Jiang et al., 2025; Zhong et al., 2024) investigate LLM agents for code generation, bug fixing, and code migration. A central factor behind their progress is inference-time iterative frameworks (Shinn et al., 2023; Lingam et al., 2025), which leverage execution feedback and self-reflection to refine candidate programs (Yang et al., 2024; Xia et al., 2025). While such methods enhance inference pipelines, they leave the base model unchanged. In contrast, our work improves the model itself through training-time optimization, strengthening the reasoning and self-correction abilities that agentic frameworks depend on.
[…]
3 Background: GRPO
Group Relative Policy Optimization (GRPO; (Shao et al., 2024)) is a variant of Proximal Policy Optimization (PPO) designed to improve the efficiency and stability of policy updates in LLM fine-tuning. Unlike PPO, which estimates advantages using a learned value function (critic), GRPO replaces the critic with an empirical baseline: the mean reward across all generations produced by the model for the same prompt. In both methods, rewards are provided by a reward model. Specifically, for each input prompt, the model samples a set of $G$ candidate responses, forming a response group. The reward model assigns a score to each response, and advantages are computed by standardizing rewards within the group: subtracting the group mean and dividing by the group standard deviation. This yields relative, normalized advantage values. As in PPO, GRPO may incorporate a penalty term to prevent the updated policy from drifting too far from the reference policy, typically enforced via a Kullback–Leibler (KL) divergence regularizer to ensure stable updates. A formal definition of the GRPO objective, along with additional details, is provided in App. B.
4 Proposed Method: Murphy
Figure 2: Overview of Murphy. Given an input prompt (q), $G$ code generations (o) are generated and evaluated using a reward function (r). Generations that do not achieve the maximum reward are revised based on executor feedback (f), combining the original prompt with the failed output, and re-prompted to generate another $G$ candidates. This iterative process continues for a fixed number of turns, with rewards from later turns propagated backward. The example illustrates the case with $G=2$, where $G$ represents the number of rollouts per prompt, and $\rho(\cdot)$ denotes the credit assignment strategy.
Extending GRPO to multi-turn interaction requires formalizing how environmental feedback informs optimization. We study a feedback-rich code generation setting where model outputs are executed and scored against test suites, yielding two feedback types: (1) quantitative (e.g., proportion of test cases passed) and (2) qualitative (e.g., error traces or failing cases). Murphy extends GRPO by introducing (i) a multi-turn rollout mechanism that conditions generation on feedback, and (ii) a credit-assignment scheme that propagates rewards from later successful turns to earlier attempts.
Multi-turn rollout: For each task prompt, the policy begins at turn 1 by generating $G_{1}$ candidate solutions, each evaluated for reward against the prompt’s test suite. Failed generations are paired with their corresponding feedback and, together with the original prompt and prior outputs, form new conditioning contexts for subsequent turns. At turn $s$, the model generates $G_{s}$ new candidates for each failed case, continuing this feedback–refinement process for a fixed number of turns. This per-prompt, iterative rollout enables progressive improvement conditioned on feedback. Unlike standard GRPO, Murphy delays advantage computation until the full rollout is complete, allowing final rewards to retroactively shape credit assignment.
Credit assignment: Once rewards for all turns are computed, Murphy propagates future rewards backward to earlier turns using a temporal credit-assignment scheme. Although early generations may achieve low initial rewards, incorporating execution feedback in later turns often leads to successful refinements. To ensure that these intermediate steps are properly credited, rewards from later successful generations are distributed to earlier ones according to their contribution to eventual success. After reward redistribution, advantages are computed for each generation, and the Murphy loss is applied to update the policy. Together, these mechanisms extend GRPO to handle multi-turn, feedback-conditioned optimization. Below, we formalize the multi-turn rollout mechanism and credit-assignment process and define the Murphy objective, extending GRPO to handle multi-turn feedback.
Notation and formalism.
We denote the model policy by $\pi_{\theta}(\cdot\mid\cdot)$ and the reference (older) policy by $\pi_{\theta_{\text{old}}}(\cdot\mid\cdot)$. Let $\mathcal{P}(Q)$ the distribution over input prompts/questions $Q$, and $O$ the output space. As described earlier, in the first turn, the model generates $G_{1}$ candidate solutions for each prompt. For generations that fail, feedback is obtained from the environment. The model is then re-prompted with the original prompt, its previous output, and the corresponding feedback to produce $G_{s}$ new generations per prompt at turn $s$. This iterative procedure naturally forms a tree structure, where the root corresponds to the original prompt, and each subsequent generation (augmented with feedback) forms a child node, with generations at turn $s+1$ linked to their parent at turn $s$ (see Fig. 2). We formalize this multi-turn rollout process as follows.
Multi-turn rollout formalism.We define a feedback-conditioned rollout tree that captures how model generations evolve across multiple turns of interaction with the environment. Let $s$ denote the turn index, $S$ the total number of turns, and $G_{s}$ the number of generations per prompt at turn $s$.
Turn 1:At the first turn ($s=1$), the model receives an input prompt $q_{(1)}$ sampled from $\mathcal{P}(Q)$ and generates $G_{1}$ candidate programs which forms a response group $\{o_{\{q_{(1)},1\}},o_{\{q_{(1)},2\}},\ldots,o_{\{q_{(1)},G_{1}\}}\}$. Note that $\{o_{\{q_{(1)},j\}}\}_{j=1}^{G_{1}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{(1)})$ . For any generation indexed by $j$, the generated code is executed against its associated test suite, producing a numerical reward $r_{\{q_{(1)},j\}}$ and environment feedback $f_{\{q_{(1)},j\}}$. In our setting, the reward represents the proportion of test cases passed, while the environment feedback consists of the specific unit tests that passed or failed, along with any corresponding error messages. These generations together form the first layer of output nodes in the rollout tree.
Turn 2:For each generation $j$ in turn 1 that fails to achieve the maximum reward (which equals 1 in our setting, since it represents the proportion of test cases passed), the corresponding feedback is appended to the original prompt and the prior output from turn $1$ to form a feedback-conditioned prompt: $q_{(2,j)}=[\ q_{(1)},\ o_{\{q_{(1)},j\}},\ f_{\{q_{(1)},j\}}\ ]$ where $\cdot\,$] denotes textual concatenation. The model is then re-invoked to generate $G_{2}$ new candidate solutions: $\{o_{\{q_{(2,j)},k\}}\}_{k=1}^{G_{2}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{(2,j)})$. Each of these generations is evaluated to obtain $(r_{\{q_{(2,j)},k\}},f_{\{q_{(2,j)},k\}})$, which denote the reward and feedback. These output generations represent refinements of their corresponding parent outputs $o_{\{q_{(1)},j\}}$ and collectively form the second layer of the rollout tree.
Turn $s$:Building on the previous turn, this procedure extends recursively to any turn $s\in\{1,\dots,S-1\}$. Note that we define:
$i_{[1:s]}=i_{1},\ \dots,\ i_{s}$
For each generation $o_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}$ that fails to achieve the maximum reward, the feedback-conditioned prompt is constructed:
$q_{(s+1,i_{[1:s]})}=[\,q_{(s,i_{[1:s-1]})},o_{\{q_{(s,i_{1:(s-1)})},i_{s}\}},f_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}]$
where $\cdot\,$] denotes textual concatenation. The model is then re-invoked to generate $G_{s+1}$ new candidate solutions $\{\,o_{\{q_{(s+1,i_{[1:s]})},k\}}\,\}_{k=1}^{G_{s+1}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{(s+1,i_{[1:s]})})$. Each generation is evaluated to obtain $(r_{\{q_{(s+1,i_{[1:s]})},k\}},\,f_{\{q_{(s+1,i_{[1:s]})},k\}})$, which denote the corresponding reward and feedback, respectively. The resulting generations, $\{o_{\{q_{(s+1,i_{[1:s]})},k\}}\,\}_{k=1}^{G_{s+1}}$ form the children nodes of the unique corresponding parent node $o_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}$. Hence, a complete path from the root (at the prompt of turn 1) to a leaf at stage $S$ (the output of the model), is expressed as:
$q_{(1)}\rightarrow o_{\{q_{(1)},i_{1}\}}\rightarrow q_{(2,i_{1})}\rightarrow o_{\{q_{(2,i_{1})},i_{2}\}}\rightarrow\cdots\rightarrow o_{\{q_{(S,i_{1},\dots,i_{S-1})},i_{S}\}}$
Note that leaf nodes at turn $S$ represent the final generations obtained after completing all refinement steps. Once rewards for all turns are computed, the rewards from these terminal nodes are propagated backward through their parent nodes according to the credit assignment strategies described below.
Credit assignment formalism.After all outputs and corresponding rewards are generated and the tree structure is constructed, we focus on assigning credit from later turns back to earlier ones. To achieve this, we employ two distinct strategies.
Max Reward Strategy (MaRS):The Max Reward Strategy is defined recursively, proceeding from the final turn back to the root. Since the final turn $S$ has no children, the rewards at this turn remain unchanged. We then consider turn s=S1s=S-1. Let $o_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}$ denote a generation at turn s=S1s=S-1 with an associated reward $r_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}$. If this generation already achieves the maximum reward, it has no children; otherwise, its children are defined as:
$C(o_{\{q_{(s,i_{1:(s-1)})},i_{s}})=\{o_{\{q_{(s+1,i_{[1:s]})},1\}},\ \ldots,\ o_{\{q_{(s+1,i_{[1:s]})},G_{S}\}}\}$
The corresponding set of rewards are defined as
$C_{r}(o_{\{q_{(s,i_{1:(s-1)})},i_{s}})=\{r_{\{q_{(s+1,i_{[1:s]})},1\}},\ \ldots,\ r_{\{q_{(s+1,i_{[1:s]})},G_{S}\}}\}$
which defaults to zero if there are no children. We then update the reward as:
$r_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}=\text{max}\Bigg(r_{\{q_{(s,i_{1:(s-1)})},i_{s}\}},\ \text{max}\Big(C_{r}(o_{\{q_{(s,i_{1:(s-1)})},i_{s}})\Big)\Bigg)$
This rule selects the greater value between a node’s own reward and the maximum reward among its descendants. Intuitively, the inner maximum represents the best improvement achievable through feedback, while the outer maximum guarantees that a node’s reward does not decrease when feedback fails to yield better performance. This formulation therefore captures the maximum attainable progress achievable from feedback at each turn. The procedure is applied recursively: rewards are first updated for all generations at turn S1S-1, then propagated backward toward the root. At each turn $s$, rewards are reassigned based on the updated values from turn $s+1$, continuing until all turns have been processed.
Mean Reward Strategy (MeRS):The Mean Reward Strategy follows the same recursive credit assignment structure as the Max Reward Strategy (MaRS), with the key difference lying in how rewards are propagated. Instead of using the maximum reward among descendants, MeRS updates each node’s reward using the mean of its child rewards, normalized over the number of stages the update has taken place over. Specifically, for a generation $o_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}$ at stage $s$, the update rule is,
$r_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}=\frac{1}{S-s+1}\Bigg(r_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}+\gamma\cdot\text{mean}\Big(C_{r}(o_{\{q_{(s,i_{1:(s-1)})}\}}\Big)\Bigg)$
where $\gamma$ is a discount factor controlling the influence of descendant rewards, and the normalization term $\frac{1}{S-s+1}$ ensures that credit assignment remains consistent across varying depths of the tree. Intuitively, this formulation captures the average improvement achieved through feedback, scaled by how many stages of updates have occurred. The normalization prevents early-stage nodes (with many descendants) from accumulating disproportionately large credit, while ensuring that terminal nodes, having no children, are not penalized. All other aspects of the recursive procedure, including the bottom-up updates from stage S1S-1 to the root, remain identical to those in MaRS.
MaRS vs MeRS:MaRS propagates the maximum descendant reward to the earlier stages, emphasizing peak performance, whereas MeRS propagates the mean reward, emphasizing stability and overall consistency. MaRS captures best-case improvement, while MeRS provides a smoother estimate of expected progress. Together, they offer complementary views of feedback-driven credit assignment.
Murphy Objective:Once the rewards are reassigned according to the chosen credit assignment strategy (MaRS or MeRS), advantages are computed similar to standard GRPO. The adjusted rewards serve as the basis for computing normalized advantages at each turn. Conditioned on a prompt $\tilde{q}$ and for $G_{s}$ generations, we normalize each reward by subtracting the mean reward and dividing by the standard deviation of the rewards obtained across the $G_{s}$ generations, yielding the normalized advantage $\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t}$. Additionally, as defined earlier, each $i$-th generation at stage $s$, $o_{\{q_{(s,i_{1:(s-1)})},i_{\}}}\in O$, corresponds to a complete output trajectory, where $o_{\{q_{(s,i_{1:(s-1)})},i_{,}t\}}$ denotes the $t$-th token and $o_{\{q_{(s,i_{1:(s-1)})},i,<t\}}$ the prefix up to (but excluding) token $t$. We denote the sequence length of the $i$-th generation by $|o_{\{q_{(s,i_{1:(s-1)})},i\}}|$. At each turn, the GRPO objective is applied using these credit-adjusted advantages. This per-turn optimization allows Murphy to incorporate feedback from later turns into earlier updates, effectively extending GRPO to a multi-turn setting. Finally, the divergence between the current and reference policies is captured by $D_{\mathrm{KL}}(\pi_{\theta}\,\|\,\pi_{{\text{ref}}})$, computed over all tokens in the generated sequences. The resulting optimization objective, which integrates credit-assigned rewards, normalized advantages, and KL regularization at each turn, defines the Murphy objective, distinguishing it from the standard GRPO formulation. The full Murphy objective is presented in Sec. 4.
Definition 1. (Murphy Objective) $\mathcal{J}_{\textsc{Murphy}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{P}(Q)}[\sum_{s,i_{1},\ldots i_{S}}\mathcal{J}_{q_{(s,i_{1:s})}}(\theta)]$ Where the per prompt objective at each turn $s$ is: $\mathcal{J}_{q_{(s,i_{1:(s-1)})}}(\theta)=\$$\mathbb{E}_{\{o_{(\tilde{q},i)}\}_{i=1}^{G_{s}}\sim\pi_{\theta_{\text{old}}}(O|\tilde{q})}\Bigg[\sum_{i=1}^{G_{s}}\frac{1}{G_{s}|o_{\{\tilde{q},i\}}|}\sum_{t=1}^{|o_{\{\tilde{q},i\}}|}\Bigg(\min\Big(R_{\theta}(\tilde{q},i,t)\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t},$ $\quad\text{clip}\left(R_{\theta}(\tilde{q},i,t),\,1-\epsilon,\,1+\epsilon\right)\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t}\Big)-\beta D_{\mathrm{KL}}(\pi_{\theta}\,\|\,\pi_{{\text{ref}}})\Bigg)\Bigg]$ with, $\tilde{q}$ $=q_{(s,i_{1:(s-1)})}$ $R_{\theta}(\tilde{q},i,t)$ $=\frac{\pi_{\theta}(o_{\{\tilde{q},i,t\}}\mid\tilde{q},o_{\{\tilde{q},i,<t\}})}{\pi_{\theta_{\text{old}}}(o_{\{\tilde{q},i,t\}}\mid\tilde{q},o_{\{\tilde{q},i,<t\}})}$ And, $\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t}$ denotes the advantage.
Note.
The design of Murphy is broadly applicable across a range of RLVR algorithms, including PPO (Schulman et al., 2017) and various extensions of GRPO (Ahmadian et al., 2024; Yu et al., 2025; Yue et al., 2025). In this work, we focus on GRPO due to its strong empirical performance in aligning LLMs (DeepSeek-AI et al., 2025; Yang et al., 2025). Extending Murphy to other RLVR variants is conceptually straightforward, as it builds on the same underlying principles. While our experiments primarily focus on code generation, where rich, verifiable feedback is readily available, the framework can naturally extend to other domains such as mathematics or logical reasoning, provided suitable forms of feedback are accessible.
4.1 Pruning Strategies in Murphy
By default, Murphy sets $G_{s}=G$ for all turns, maintaining a fixed number of generations per prompt throughout the multi-turn process. However, this multi-turn setup introduces significant computational cost. In the worst case, when success is achieved only at the final turn $S$, the number of generations per prompt can grow exponentially to $G^{S}$, resulting in a large search tree and substantial memory overhead. This makes Murphycomputationally expensive to optimize. While system-level optimizations such as vLLM with paged attention and KV caching (Kwon et al., 2023) make the generation process relatively efficient, the optimization phase remains costly. Large-scale rollouts can quickly exhaust GPU memory, and batching schemes that treat each generated sample as an independent training batch tend to be prohibitively slow. To address these challenges, we introduce two pruning strategies that reduce the number of rollouts at each turn, thereby making Murphy computationally tractable without compromising performance. We describe these strategies in detail below.
Intra-Group Pruning (IntraP):In IntraP, pruning operates recursively: starting from the children of stage S1S-1, each pruning step is followed by reward propagation as described in Sec. 4, and the process continues backward until reaching the root. This ordering—pruning before reward reassignment—ensures that only informative trajectories are retained for credit propagation across turns. At each turn $s$, given a pruning budget $b$, we retain only the $b$ trajectories conditioned on the prompt whose rewards contribute most to the total reward variance within that group. The remaining trajectories, along with all their descendants, are discarded, and the process proceeds recursively until the root is reached. This approach is inspired by Xu et al. (2025a), who demonstrate that retaining trajectories with the highest reward variance within a group can reduce optimization cost while maintaining performance comparable to GRPO. We extend this principle to the multi-turn setting of Murphy.
Inter-Group Pruning (InterP):In InterP, the goal is to prune entire groups of children corresponding to generations conditioned on a given prompt. Specifically, each generation at a turn has a group of children, and we decide how many of these groups to retain based on a pruning budget $b$. To rank the groups, we assign a score inspired by the UCB sampling strategy (Auer et al., 2002), defined as $\alpha_{1}\mu+\alpha_{2}\sigma$, where $\mu$ and $\sigma$ denote the mean and standard deviation of the rewards within each group of children. The groups are ranked in descending order of this score, and only the top $b$ groups are retained, with $b$ determining the available computational budget. The remaining groups and their descendants are discarded. Similar to IntraP, this process is applied recursively and proceeds backward from the latter turns to the earlier ones. At each level, pruning is performed first, followed by credit assignment. In our experiments, we set $\alpha_{1}=0$ and $\alpha_{2}=1$, which effectively prioritizes groups exhibiting higher reward variance. Intuitively, a higher standard deviation not only captures greater variability within a group but also indicates that the model is uncertain about that group’s performance. Such groups likely contain some generations with near-maximum rewards, making them valuable targets for further optimization. In contrast, groups where all rewards are uniformly high (near $1$) or uniformly low (near 0) provide limited learning signal, as the model either has already mastered or completely failed the underlying behavior. High-variance groups, therefore, represent the most informative regions for continued improvement.
5 Experiments
In this section, we first provide an overview of the models and datasets used in our experiments, followed by a detailed description of the setup and results, including ablation studies. We provide implementation details in App. D.
Models. We evaluate our framework using two open-source model families: Qwen3 (1.7B, 4B) (Yang et al., 2025) and OLMo2 (OLMo-2-1124-7B-Instruct) (OLMo et al., 2025). These models differ in both architecture and scale, enabling us to assess the generality of our method across diverse settings.
Datasets, Metrics, and Evaluation. We detail the datasets, metrics, and evaluation procedures employed in our experiments below.
Training Dataset: We fine-tune all models on 1,000 samples randomly drawn from the KodCode dataset (Xu et al., 2025b). This dataset was chosen due to its minimal overlap with evaluation benchmarks [^1] Refer to Section 3.2 of KodCode (Xu et al., 2025b) for details on contamination analysis.. We further study the impact of training data size in Appendix D.4.2.
Evaluation Datasets. We evaluate trained models on a suite of programming and reasoning benchmarks, including HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and BigCodeBench-Hard (Zhuo et al., 2025). For BigCodeBench-Hard, which does not include visible unit tests, we randomly sample two test cases from the full test suite to construct visible tests.
Metrics and Evaluation Protocol. Reflexion Shinn et al. (2023) is a widely adopted inference-time iterative framework that enhances reasoning by refining incorrect outputs through feedback and self-generated reflections (see App. C for details). To assess reasoning refinement and self-correction, we integrate all models into the Reflexion framework and report pass@1 under two settings: (i) Single iteration, equivalent to standard input–output prompting, and (ii) Three iterations, where feedback from visible test cases is incorporated into subsequent prompts. The agent terminates once all visible tests pass or when the maximum iteration limit is reached. Final solutions are evaluated on hidden test cases, and the resulting pass@1 is reported. Each experiment is repeated three times, and we report the mean and standard deviation of pass@1 across runs.
5.1 Murphy Experiments
We evaluate three models, Qwen3 (1.7B, 4B) (Yang et al., 2025) and OLMo-2-1124-7B-Instruct (OLMo et al., 2025), each fine-tuned on 1,000 samples randomly drawn from the KodCode dataset (Xu et al., 2025b). Evaluation is conducted on the benchmark datasets described in Sec. 5 using the Reflexion framework. We report pass@1 performance under both single- and multi-iteration settings, as summarized in Tab. 1. Additional results comparing with a naive multi-turn extension are provided in Appendix D.4.1, and experiments with higher multi-turn training are included in Appendix D.4.3. For all our tables, best results are bold and second best are underlined. Results are reported as pass@1 accuracy (% mean $\pm$ stdev).
Reflexion: Single-iteration setting. (Iter-1 in Tab. 1) Models trained with the GRPO objective consistently outperform their base counterparts, achieving notable gains; for instance, a $\sim 3\%$ improvement on HumanEval for OLMo-2-1124-7B-Instruct. Murphy achieves competitive or superior performance than GRPO in this setting.
Reflexion: Multi-iteration setting. (Iter-3 in Tab. 1) We repeat the experiments with three iterations in the Reflexion framework to assess self-correction and reasoning-refinement capabilities. Increasing the number of iterations leads to consistent performance improvements across all models and benchmarks. Models trained with Murphy surpass both GRPO-trained and base models, achieving gains of up to $8\%$ over GRPO. These results demonstrate the effectiveness of multi-turn reflective optimization in enhancing reasoning refinement and self-correction.
Model Rollouts HumanEval (%) MBPP (%) BigCodeBench (%)
Iter-1 Iter-3 Iter-1 Iter-3 Iter-1 Iter-3
Qwen3-1.7B
   Base 74.19 $\pm$ 1.95 80.07 $\pm$ 1.27 43.47 $\pm$ 0.64 53.93 $\pm$ 3.52 7.20 $\pm$ 2.73 18.24 $\pm$ 1.79
   GRPO 72 77.42 $\pm$ 0.59 82.11 $\pm$ 1.76 45.93 $\pm$ 1.60 57.20 $\pm$ 1.40 10.58 $\pm$ 1.95 20.49 $\pm$ 2.17
   Murphy- MaRS 72 79.67 $\pm$ 3.01 86.58 $\pm$ 1.06 44.73 $\pm$ 0.50 62.00 $\pm$ 1.91 6.98 $\pm$ 3.72 20.25 $\pm$ 3.07
Qwen3-4B
   Base 90.04 $\pm$ 3.13 93.49 $\pm$ 0.93 52.13 $\pm$ 0.42 70.93 $\pm$ 1.01 17.56 $\pm$ 1.78 36.03 $\pm$ 3.05
   GRPO 72 88.61 $\pm$ 0.93 94.71 $\pm$ 0.93 51.73 $\pm$ 0.23 72.87 $\pm$ 0.90 20.95 $\pm$ 0.67 39.64 $\pm$ 2.73
   Murphy- MaRS 72 92.48 $\pm$ 0.61 95.73 $\pm$ 0.31 53.33 $\pm$ 1.15 73.33 $\pm$ 1.31 22.52 $\pm$ 2.47 41.44 $\pm$ 1.09
OLMo-2-1124-7B-Instruct
   Base 37.20 $\pm$ 0.86 46.04 $\pm$ 0.43 19.90 $\pm$ 0.42 29.60 $\pm$ 0.28 1.35 $\pm$ 0.00 3.72 $\pm$ 0.48
   GRPO 72 41.26 $\pm$ 0.35 43.70 $\pm$ 0.93 32.87 $\pm$ 0.23 35.80 $\pm$ 1.00 0.68 $\pm$ 0.00 2.70 $\pm$ 0.68
   Murphy- MaRS 72 45.53 $\pm$ 0.70 52.24 $\pm$ 1.96 29.33 $\pm$ 0.90 39.67 $\pm$ 1.29 1.80 $\pm$ 0.39 3.38 $\pm$ 1.17
Table 1: Performance of Qwen3-1.7B, Qwen3-4B, and OLMo-2-1124-7B-Instruct variants across evaluation benchmarks. The Rollouts column indicates the total number of generations across both stages. Murphy achieves competitive (single-turn; iter-1) or superior results (multi-turn; iter-3), with gains of up to $\mathbf{8\%}$} over GRPO under an equivalent compute budget.
Model & Strategy Rollouts HumanEval (%) MBPP (%) BigCodeBench (%)
Iter-1 Iter-3 Iter-1 Iter-3 Iter-1 Iter-3
Qwen3-1.7B
   MaRS 72 79.67 $\pm$ 3.01 86.58 $\pm$ 1.06 44.73 $\pm$ 0.50 62.00 $\pm$ 1.91 6.98 $\pm$ 3.72 20.25 $\pm$ 3.07
   MeRS ($\gamma=0.9$) 72 78.66 $\pm$ 2.11 84.76 $\pm$ 1.22 46.53 $\pm$ 0.81 60.53 $\pm$ 1.79 10.81 $\pm$ 0.00 22.52 $\pm$ 5.67
   MeRS ($\gamma=1$) 72 78.46 $\pm$ 0.70 85.57 $\pm$ 0.35 45.07 $\pm$ 1.10 60.93 $\pm$ 1.29 9.46 $\pm$ 2.44 20.27 $\pm$ 1.35
OLMo-2-1124-7B-Instruct
   MaRS 72 45.53 $\pm$ 0.70 52.24 $\pm$ 1.96 29.33 $\pm$ 0.90 39.67 $\pm$ 1.29 1.80 $\pm$ 0.39 3.38 $\pm$ 1.17
   MeRS ($\gamma=1)$ 72 37.40 $\pm$ 0.35 41.87 $\pm$ 1.27 27.87 $\pm$ 0.12 34.40 $\pm$ 2.40 0.90 $\pm$ 0.39 1.13 $\pm$ 0.39
Table 2: Ablation study comparing Max Reward (MaRS) vs Mean Reward (MeRS) propagation strategies across Qwen3-1.7B and OLMo-2-1124-7B-Instruct. MaRS consistently outperforms MeRS in both single (Iter-1) and multi-turn (Iter-3).
5.2 Ablation 1: MaRS vs. MeRS
As described in Sec. 4, we compare two reward propagation strategies: MaRS and MeRS. We train Qwen3-1.7B and OLMo-2-1124-7B-Instruct on 1,000 KodCode samples under each strategy and report results in Tab. 2. Across both models and multiple Reflexion iterations, MaRS consistently matches or surpasses MeRS, independent of the discount factor $\gamma$. The key difference lies in handling non-binary rewards: MeRS averages rewards across generations, diluting the learning signal when few outputs score highly, whereas MaRS propagates the strongest outcome, allowing rare but valuable high-reward trajectories to dominate the update. This makes MaRS particularly effective in multi-turn settings with sparse rewards, while the gap between the two strategies diminishes in binary-reward tasks.
5.3 Ablation 2: IntraP vs. InterP
In Subsec. 4.1, we introduced two pruning strategies: Intra-Group (IntraP) and Inter-Group (InterP) pruning. Both variants prune the rollouts tree, thereby reducing the total number of gradient updates (denoted as Updates in Tab. 3). Here, we compare the effectiveness of these strategies and observe that InterP consistently matches or surpasses the performance of Murphy without pruning. This indicates that pruning across groups offers a more robust mechanism for eliminating redundant rollouts while retaining high-quality candidates. The ablation study is conducted using the Qwen3-1.7B model, and the results are reported in Tab. 3.
Model & Strategy Updates HumanEval (%) MBPP (%) BigCodeBench (%)
Iter-1 Iter-3 Iter-1 Iter-3 Iter-1 Iter-3
Murphy (MaRS) 72 79.67 $\pm$ 3.01 86.58 $\pm$ 1.06 44.73 $\pm$ 0.50 62.00 $\pm$ 1.91 6.98 $\pm$ 3.72 20.25 $\pm$ 3.07
Murphy (MaRS) – IntraP 36 82.34 $\pm$ 1.03 84.28 $\pm$ 2.01 45.73 $\pm$ 0.50 59.87 $\pm$ 1.70 11.03 $\pm$ 2.81 21.17 $\pm$ 2.07
Murphy (MaRS) – InterP 40 77.43 $\pm$ 2.20 86.17 $\pm$ 0.70 44.53 $\pm$ 0.90 61.20 $\pm$ 1.39 10.58 $\pm$ 2.56 24.54 $\pm$ 2.82
Table 3: Comparison of Murphy and its pruned variants on evaluation benchmarks using Qwen3-1.7B. All variants generate 72 rollouts per query. The Updates column denotes the total number of gradient steps per query. Pruned variants (IntraP, InterP) maintain competitive performance relative to the non-pruned version with significant computation reduction.
6 Conclusion
We introduced Murphy, a multi-turn reflective reinforcement learning framework that extends RLVR algorithms by incorporating iterative self-correction through both quantitative and qualitative feedback. By grounding optimization in intermediate feedback signals and propagating rewards across refinement turns, Murphy consistently enhances reasoning and code generation performance over GRPO, particularly in multi-iteration settings where feedback-driven refinement is crucial. These findings underscore the value of integrating structured feedback directly into the optimization process. Future directions include developing adaptive rollout selection strategies, extending Murphy to broader reasoning domains, and coupling it with inference-time search to further strengthen both training- and test-time reasoning.
7 Limitations
While effective, Murphy’s multi-turn design increases computational cost; pruning partially mitigates this but it remains more resource-intensive than single-stage baselines. Our experiments are limited to structured feedback in code generation, leaving open questions about generalization to noisier feedback, deeper refinement chains, and broader notions of agentic performance such as robustness and alignment with developer intent.
8 Acknowledgments
We acknowledge that Fig. 2 incorporates icons sourced from Flaticon. Full creator attributions are provided in App. G.
Appendix
Appendix A Extended Related Work
LLM Agents for Software Development.
Recent works (Jiang et al., 2025; Zhong et al., 2024) have explored LLM agents for programming tasks such as code generation, bug fixing, and code migration. A key driver of progress in these domains has been inference-time iterative frameworks (Shinn et al., 2023; Lingam et al., 2025), which leverage execution feedback to generate self-reflections for refining candidate programs (Yang et al., 2024; Xia et al., 2025). While these approaches underscore the value of iterative feedback and scaffolding, they primarily enhance the inference pipeline rather than the underlying model. Our work takes a complementary direction: we improve the reasoning and self-correction abilities of LLMs themselves through training-time optimization, thereby strengthening the base models that agentic frameworks depend on.
Reinforcement Learning with Verifiable Rewards for LLM Reasoning.
[…]
Appendix B GRPO: Objective and Additional Details
Notation.
We denote the model policy by $\pi_{\theta}(\cdot\mid\cdot)$ and the reference (older) policy by $\pi_{\theta_{\text{old}}}(\cdot\mid\cdot)$. Let $G$ be the number of generations per prompt, $\mathcal{P}(Q)$ the distribution over input prompts/questions $Q$, and $O$ the output space. For a given prompt $q\sim\mathcal{P}(Q)$, the reference policy produces a set of $G$ responses, forming a response group $\{o_{q,1},o_{q,2},\ldots,o_{q,G}\}$. Each generation $o_{q,i}\in O$ corresponds to a full output trajectory, where $o_{q,i,t}$ denotes the $t$-th token and $o_{q,i,<t}$ the prefix up to (but excluding) token $t$. We write $|o_{q,i}|$ for the sequence length of the $i$-th generation. For a given prompt $q$, the reward model assigns a scalar score to each response in the group, yielding $\mathbf{r}_{q}=\{r_{q,1},r_{q,2},\ldots,r_{q,G}\}$. Moreover, for prompt $q$, the advantage associated with the $t$-th token of the $i$-th generation is defined as $\hat{A}_{q,i,t}=(r_{q,i}-\mu(\mathbf{r}_{q}))/\sigma(\mathbf{r}_{q})$, where $\mu(\mathbf{r}_{q})$ and $\sigma(\mathbf{r}_{q})$ denote the mean and standard deviation of the group rewards, respectively. Finally, $D_{\mathrm{KL}}(\pi_{\theta}||\pi_{\theta_{\text{old}}})$ denotes the KL divergence between the current and reference policies, computed over all tokens in the generated sequences. The GRPO training objective is presented in App. B.
Definition 2. (GRPO Objective) $\mathcal{J}(\theta)=\$$\mathbb{E}_{q\sim\mathcal{P}(Q),\ \{o_{(q,i)}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}\Big[\frac{1}{G}\sum_{i=1}^{G}$ $\frac{1}{|o_{q,i}|}\sum_{t=1}^{|o_{q,i}|}\min\Big(R_{\theta}(q,i,t)\hat{A}_{q,i,t},$ $\text{clip}\Big(R_{\theta}(q,i,t),\,1-\epsilon,\,1+\epsilon\Big)\hat{A}_{q,i,t}\Big)\Big]$ $-\beta D_{\text{KL}}(\pi_{\theta}\,\|\,\pi_{\text{ref}})$ Where, $R_{\theta}(q,i,t)$ $=\frac{\pi_{\theta}(o_{q,i,t}\mid q,o_{q,i,<t})}{\pi_{\theta_{\text{old}}}(o_{q,i,t}\mid q,o_{q,i,<t})}$ $\hat{A}_{q,i,t}$ $=\frac{(r_{q,i}-\mu(\mathbf{r}_{q}))}{\sigma(\mathbf{r}_{q})}$
Appendix C Reflexion
Reflexion (Shinn et al., 2023) is an inference-time iterative framework designed to improve reasoning through repeated interaction with feedback from an external environment. It employs three agents: an actor ($M_{a}$), an evaluator ($M_{e}$), and a self-reflection module ($M_{sr}$), which operate cyclically until a termination condition is met. For code generation, the process proceeds as follows:
  1. 1.
    Actor step: The actor $M_{a}$ receives an input and generates an output (e.g., a code snippet).
  2. 2.
    Evaluation step: The evaluator $M_{e}$ scores the output (e.g., the percentage of unit tests passed).
  3. 3.
    Self-reflection step: If the score is insufficient, the self-reflection module $M_{sr}$ diagnoses the issue, proposes a fix, and appends both the failed output and the suggested correction to the input context. The updated input is then fed back to the actor, and the cycle repeats until either the task succeeds or a maximum number of iterations is reached.
In most implementations, the actor and the self-reflection module are instantiated by the same underlying language model. The self-reflection stage thus corresponds to the model reasoning over its own prior outputs augmented with feedback from the evaluator (or executor) and the previous input, output pairs, to generate improved responses in subsequent iterations.
Appendix D Implementation Details
We implement our framework on top of TRL (von Werra et al., 2020), which provides efficient distributed training and a modular implementation of GRPO. We integrate TRL with vLLM for fast inference and large-scale rollout execution, enabling scalable multi-turn training in our experiments. Prompts used to train Murphy are listed in App. E. All experiments use publicly available datasets. The base models (Qwen3, OLMo) are available for research use under the Apache 2.0 license.
D.1 Model Size and Compute Budget
All experiments were conducted on $8\times$ NVIDIA H100 GPUs. Our implementation builds on HuggingFace’s TRL [^2] https://huggingface.co/docs/trl/en/index. For efficiency, 2 GPUs were allocated for inference via vLLM, while the remaining 6 GPUs handled model updates. Training Qwen3-1.7B with Murphy on 1,000 KodCode samples took approximately 1.5 hours, Qwen3-4B took 4 hours, and OLMo-2-7B-Instruct required 10 - 13 hours. Checkpoints were saved every 50 steps, and for all baselines, we selected the checkpoint corresponding to one epoch.
D.2 Hyperparameters
We set the KL regularization factor $\beta=0.04$, learning rate to $10^{-6}$, and weight decay to $0.1$ for both GRPO and Murphy  variants. Unless stated otherwise, the number of stages in Murphy is set to 2. For GRPO and the first stage of Murphy, we use 8 rollouts per prompt, while the second stage uses up to 8 rollouts per prompt (resulting in a maximum of 64 rollouts). To ensure a fair computational comparison, we also train GRPO with 72 rollouts. Following Yang et al., we set the temperature to 0.6 and top-$p$ to 0.95 for all experiments.
D.3 Package Parameters
We use the Reflexion (Shinn et al., 2023) framework to evaluate all trained models. The number of iterations is swept over $\{1,3\}$, and max-tokens is set to each model’s maximum generation length. Models are hosted via vLLM. Since all models fit on a single H100 GPU, we set data-parallel-size to 8 and enable prefix caching to accelerate evaluation. We use the following commands to install the appropriate packages:
pip install uv && \
uv pip install trl==0.19.1 && \
uv pip install gunicorn==20.1.0 && \
uv pip install fastapi==0.115.12
uv pip install uvicorn==0.34.2 && \
uv pip install aiohttp==3.11.18
uv pip install astunparse==1.6.3
uv pip install jsonlines tenacity && \
uv pip install vllm==0.8.5.post1
D.4 Additional Experiments
D.4.1 Ablation: Comparison with a Naive Multi-Turn Extension
A straightforward way to adapt GRPO for multi-turn training is to extend failed rollouts by appending feedback from previous turns without introducing additional fan-out. Rewards and advantages are then computed on the final turn, and the GRPO objective is applied. We compare this baseline against Murphy in Tab. 4. The number of turns is set to 2. The results underscore the importance of Murphy’s structured credit assignment and staged fan-out, both of which contribute to its superior multi-turn optimization performance.
D.4.2 Ablation: Effect of Training Dataset Size
To examine the effect of training dataset size, we construct three nested subsets of KodCode ($2\subset$ $3\subset$ 44K) and train Qwen3-1.7B using GRPO (72 rollouts) and Murphy. Results are summarized in Tab. 5. While performance does not increase monotonically with dataset size, Murphy consistently demonstrates superior multi-turn robustness and maintains an average improvement of up to $\sim 9\%$ over GRPO across all dataset scales.
D.4.3 Ablation: Effect of Pruning in Multi-Turn Murphy Training
We noted in the main paper that increasing the number of turns in Murphy can lead to exponential growth in computational cost. To mitigate this, we design and evaluate two pruning strategies. In this experiment, we extend our setup to a 3-turn setting and study the effect of pruning on performance. Results in Tab. 6 show that the pruned variant achieves competitive or even superior performance compared to the non-pruned counterpart.
Model Rollouts HumanEval (%) MBPP (%) BigCodeBench (%)
Iter-1 Iter-3 Iter-1 Iter-3 Iter-1 Iter-3
Qwen3-1.7B
   Base 74.19 $\pm$ 1.95 80.07 $\pm$ 1.27 43.47 $\pm$ 0.64 53.93 $\pm$ 3.52 7.20 $\pm$ 2.73 18.24 $\pm$ 1.79
   GRPO 72 77.42 $\pm$ 0.59 82.11 $\pm$ 1.76 45.93 $\pm$ 1.60 57.20 $\pm$ 1.40 10.58 $\pm$ 1.95 20.49 $\pm$ 2.17
   Murphy (Simple) 144 70.33 $\pm$ 0.93 82.11 $\pm$ 0.35 46.00 $\pm$ 0.40 58.53 $\pm$ 1.03 13.29 $\pm$ 1.41 18.24 $\pm$ 1.17
   Murphy- MaRS 72 79.67 $\pm$ 3.01 86.58 $\pm$ 1.06 44.73 $\pm$ 0.50 62.00 $\pm$ 1.91 6.98 $\pm$ 3.72 20.25 $\pm$ 3.07
OLMo-2-1124-7B-Instruct
   Base 37.20 $\pm$ 0.86 46.04 $\pm$ 0.43 19.90 $\pm$ 0.42 29.60 $\pm$ 0.28 1.35 $\pm$ 0.00 3.72 $\pm$ 0.48
   GRPO 72 41.26 $\pm$ 0.35 43.70 $\pm$ 0.93 32.87 $\pm$ 0.23 35.80 $\pm$ 1.00 0.68 $\pm$ 0.00 2.70 $\pm$ 0.68
   Murphy (Simple) 144 39.02 $\pm$ 0.00 41.87 $\pm$ 0.93 28.33 $\pm$ 0.31 33.60 $\pm$ 0.69 1.13 $\pm$ 1.41 2.25 $\pm$ 0.39
   Murphy- MaRS 72 45.53 $\pm$ 0.70 52.24 $\pm$ 1.96 29.33 $\pm$ 0.90 39.67 $\pm$ 1.29 1.80 $\pm$ 0.39 3.38 $\pm$ 1.17
Table 4: Performance of Qwen3-1.7B and OLMo-2-1124-7B-Instruct variants on HumanEval, MBPP, and BigcodeBench benchmarks, reported as pass@1 accuracy (% mean $\pm$ stdev over 3 independent runs). Best results are bold, second-best are underlinedMurphy- MaRS outperforms Murphy  (Simple) on average.
Qwen3-1.7B Dataset HumanEval HumanEval MBPP MBPP BigcodeBench BigcodeBench
Iter-1 Iter-3 Iter-1 Iter-3 Iter-1 Iter-3
Murphy- MaRS 2K $\mathbf{81.65\%\pm 2.19}$ $\mathbf{84.15\%\pm 2.05}$ $\mathbf{46.93\%\pm 0.81}$ $\mathbf{60.47\%\pm 2.66}$ $\mathbf{2.70\%\pm 0.68}$ $\mathbf{11.04\%\pm 1.70}$
GRPO 2K $78.46\%\pm 2.46$ $83.54\%\pm 1.06$ $46.13\%\pm 0.70$ $57.67\%\pm 0.61$ $2.48\%\pm 0.78$ $8.11\%\pm 1.17$
Murphy- MaRS 3K $77.30\%\pm 3.55$ $\mathbf{84.55\%\pm 1.96}$ $45.53\%\pm 0.12$ $\mathbf{63.60\%\pm 0.72}$ $\mathbf{3.83\%\pm 0.78}$ $\mathbf{10.36\%\pm 1.41}$
GRPO 3K $\mathbf{79.07\%\pm 2.54}$ $80.49\%\pm 1.61$ $\mathbf{47.40\%\pm 1.20}$ $54.87\%\pm 1.68$ $3.60\%\pm 1.03$ $6.08\%\pm 0.68$
Murphy- MaRS 4K $79.27\%\pm 2.26$ $\mathbf{89.63\%\pm 1.22}$ $45.60\%\pm 1.31$ $\mathbf{62.80\%\pm 0.92}$ $\mathbf{3.60\%\pm 0.39}$ $\mathbf{8.78\%\pm 2.70}$
GRPO 4K $\mathbf{79.88\%\pm 2.66}$ $83.33\%\pm 0.70$ $\mathbf{47.27\%\pm 0.46}$ $57.40\%\pm 1.83$ $2.70\%\pm 1.17$ $6.98\%\pm 1.41$
Table 5: Ablation: average $\pm$ stdev (percent) for Murphy vs GRPO on KodCode (2K/3K/4K). Murphy MaRS shows gains of up to $\sim 9\%$ over compute equivalent GRPO.
Model HumanEval MBPP BigCodeBench
Iter-1 Iter-3 Iter-1 Iter-3 Iter-1 Iter-3
GRPO $77.24\pm 2.14$ $83.13\pm 1.96$ $44.07\pm 0.31$ $53.93\pm 0.61$ $1.13\pm 0.39$ $6.53\pm 0.78$
Murphy- MaRS, InterP $78.66\pm 2.66$ $84.15\pm 1.83$ $45.53\pm 1.17$ $61.20\pm 0.53$ $2.70\pm 0.68$ $7.43\pm 1.79$
Murphy- MaRS $79.27\pm 2.11$ $83.94\pm 1.41$ $44.80\pm 1.11$ $62.00\pm 1.40$ $1.35\pm 0.68$ $5.86\pm 0.39$
Table 6: Ablation study comparing pruning versus non-pruning strategies for Qwen3-1.7B trained with Murphy for 3 turns on the KodCode dataset. Reported numbers indicate pass@1 (%) over three independent runs. Pruned variant achieves competitive or superior performance compared to non-pruned Murphy.
Appendix E Prompt Examples
To train the Murphy objective, we employ the following prompts. The system prompt is used at each dialogue turn, and the feedback prompts are applied during every feedback turn.
System Prompt You are a Python coding AI agent. When given a Python function signature and docstring, you must provide a complete Python solution following this exact format: 1. Reasoning Phase: Use <think>...</think> tags to contain your complete thought process. Break down the problem requirements step by step Identify key constraints, edge cases, and potential pitfalls Plan your algorithm and data structures Walk through examples to validate your approach Explain your logic thoroughly as if working on scratch paper 2. Implementation Phase: Use <output>...</output> tags to contain your final code. Include the complete function with the original signature Ensure your code directly implements the approach from your thinking Write clean, readable code with appropriate comments if needed Your solution must be complete, correct, and handle all specified requirements.
Feedback Header You have previously attempted this problem {num-attempts} time(s).
Feedback Body Previous Attempt #{num-attempts} Analysis: Your earlier thought process and implementation: Let me think step by step. <think> {code-generated} Test Case Results: - Passed: {passed-tests}/{total-tests} test cases - Failed: {failed-tests}/{total-tests} test cases Detailed Feedback: {feedback-string} Note: For failed test cases, your code’s output is shown followed by a ‘#‘ symbol indicating the failure.
Feedback Footer. Instructions for Your Next Attempt: 1. Failure Analysis Phase: (a) Carefully examine each failed test case to understand exactly what went wrong (b) Identify the specific lines of code or logic that caused the failures (c) Look for patterns across multiple failed cases (e.g., all involve negative numbers, empty inputs, etc.) (d) Determine if failures stem from algorithmic errors, edge case handling, or implementation bugs 2. Solution Refinement Phase: (a) Build upon any correct aspects of your previous solution that passed tests (b) Redesign the problematic parts of your algorithm to handle the failed cases (c) Ensure your new approach covers edge cases that weren’t properly handled before (d) Consider additional edge cases that might not be in the test suite but could break your solution 3. Implementation Phase: (a) Write your improved solution that directly addresses the identified failure points (b) Test your logic mentally against the failed cases to verify it would now pass (c) Ensure your solution maintains correctness for previously passing test cases Use the concrete feedback from your {num-attempts} previous attempt(s) to create a more robust and accurate solution.
Appendix F Potential Risks
While Murphy introduces some new dynamics through iterative self-correction and reflective optimization, the associated risks appear modest overall. The main considerations involve ensuring that feedback loops remain InterPretable and that reward signals do not inadvertently reinforce narrow or heuristic reasoning. There is also some potential for subtle reward hacking, where the model optimizes for easily verifiable but shallow improvements, or for mild distributional drift if reflective heuristics fail to generalize beyond training contexts. Nonetheless, because Murphy still relies on verifiable rewards and bounded reflection, these risks are relatively contained and can be mitigated through careful evaluation design, human oversight, and robust validation across diverse task settings.
Appendix G Icon Attributions