### v2

# Murphy: Multi-Turn GRPO for Self-Correcting Code Generation

Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has emerged as a powerful framework for enhancing the reasoning capabilities of large language models (LLMs). However, existing approaches such as Group Relative Policy Optimization (GRPO) and its variants, while effective on reasoning benchmarks, struggle with agentic tasks that require iterative decision-making. We introduce Murphy, a multi-turn RLVR framework that incorporates execution feedback directly into training, extending GRPO to optimize over multi-turn trajectories where models iteratively refine solutions. Murphy combines a feedback-conditioned rollout tree with trajectory-level credit assignment, and uses pruning to reduce the cost of multi-turn optimization. Evaluations on code generation benchmarks with two model families show that Murphy consistently improves multi-iteration performance, achieving up to an $8\%$ absolute gain in pass@1 over compute-matched GRPO baselines, and outperforming the prior leading method that incorporates multi-turn execution feedback.
Machine Learning, ICML
1 Introduction
“The road to wisdom? Well, it’s plain and simple to express: err and err and err again, but less and less and less.” —Piet Hein
Qwen3-1.7BQwen3-4BOLMo-2-7B-Ins011223344556677882.82.81.41.43.23.26.76.72.92.97.47.4ModelsPercentage (%)($\%$)GRPOMurphy
Figure 1: Percentage change in coding problems solved by models trained with Murphy and GRPO over the base model across three models and datasets. For Qwen3-4B, we additionally include results on the Aider Polyglot benchmark. Murphy-trained models solve up to $4.2\%$ more problems than GRPO. See Tab. 1 and Subsec. 5.1 for details.
A growing body of work explores large language models (LLMs) as software engineering agents that interact with their environment through code execution and feedback (Shinn et al., 2023; Lingam et al., 2025; Miserendino et al., 2025). Rather than producing a single static response, these systems operate within agentic scaffolds (Yao et al., 2023; Zhou et al., 2024) that guide iterative reasoning and allow LLMs to act, observe, and improve over multiple rounds of interaction. For example, in a typical coding agentic scaffold (Shinn et al., 2023; Lingam et al., 2025), the agent generates a solution by performing a single or series of actions, and executes it for evaluation, often through unit tests or other automated checks. When execution fails, the agent receives feedback such as error messages, stack traces, or failing test cases, and is re-prompted with the original task, its previous attempt, and the new feedback. This process continues for multiple turns until the model succeeds or reaches a fixed iteration limit. These systems highlight the growing ability of LLMs to reason and adapt through environmental feedback at inference time. In code generation tasks (Jiang et al., 2025), such feedback naturally arises from execution logs, compiler errors, or test results. However, these methods (Shinn et al., 2023; Lingam et al., 2025) remain fundamentally training-free: they improve model behavior through structured inference via textual feedback rather than through parameter updates.
On the other hand, training methodologies such as Reinforcement Learning with Verifiable Rewards (RLVR) have enabled a new generation of language models (OpenAI et al., 2024; DeepSeek-AI et al., 2025; Team et al., 2025; Yang et al., 2025) to exhibit strong reasoning capabilities across mathematics, coding, and general problem-solving. Recent RLVR algorithms, including Group Relative Policy Optimization (GRPO) (Shao et al., 2024) and its extensions (Yue et al., 2025; Yu et al., 2025; Xu et al., 2025a), have become dominant approaches for post-training LLMs on verifiable reasoning tasks. However, GRPO and its extensions are fundamentally designed for a single-turn setting: they optimize model behavior using an isolated prompt–response–reward tuple, with no notion of multi-turn interaction defined in their objective. Here, a turn denotes one cycle in which the model receives a prompt (which includes feedback from its previous response), reasons based on the prompt, and generates a revised output. Taken together, these limitations underscore a methodological gap: while inference-time agentic frameworks exploit iterative feedback to refine reasoning, GRPO optimizes solely from terminal rewards on a given prompt, without leveraging intermediate environmental feedback. This motivates the following key research question.
Key Research Question: How can RLVR algorithms be extended to integrate iterative feedback, and to what extent does this improve reasoning and performance within agentic frameworks?
To address this question, we propose Murphy, a novel RLVR algorithm that extends GRPO to a multi-turn setting by conditioning optimization on intermediate environmental feedback. Extending GRPO beyond a single turn is non-trivial: it requires defining how rewards obtained in later turns should be propagated backward to earlier attempts, so that intermediate reasoning and output turns that initially failed but ultimately led to success through feedback receive appropriate credit. At the first turn, Murphy generates $G_{1}$ generations per prompt and computes group-based rewards as in GRPO. In subsequent turns, generations that fail to achieve the maximum reward, e.g., those failing unit tests or producing incorrect outputs, are refined using signals from the environment such as executor logs or test results. This feedback is appended to the original prompt, and the model is re-prompted to generate a new batch of $G_{s}$ generations (where $s$ denotes the turn) conditioned on the combined context (prompt, previous output, and feedback), repeating this process for a fixed number of turns. Rewards from successful final-turn rollouts are then propagated backward to earlier turns using Murphy’s credit-assignment criterion, allowing partially correct but improving attempts to receive credit. To manage the computational cost of multi-turn updates, Murphy employs pruning strategies that retain only promising trajectories while bounding total gradient updates per turn. In summary, our main contributions are:
[…]
This work focuses on code generation tasks where execution provides structured, verifiable feedback, enabling us to measure feedback-driven refinement using agentic scaffolds at evaluation time.
2 Related Work
LLM Agents for Software Development.
Recent studies (Jiang et al., 2025; Zhong et al., 2024) investigate LLM agents for code generation, bug fixing, and code migration. A central factor behind their progress is inference-time iterative frameworks (Shinn et al., 2023; Lingam et al., 2025), which leverage execution feedback and self-reflection to refine candidate programs (Yang et al., 2024; Xia et al., 2025). While such methods enhance inference pipelines, they leave the base model unchanged. In contrast, our work improves the model itself through training-time optimization, strengthening the reasoning and self-correction abilities that agentic frameworks depend on.
[…]
3 Background: GRPO
Group Relative Policy Optimization (GRPO; (Shao et al., 2024)) is a variant of Proximal Policy Optimization (PPO) designed to improve the efficiency and stability of policy updates in LLM fine-tuning. Unlike PPO, which estimates advantages using a learned value function (critic), GRPO replaces the critic with an empirical baseline: the mean reward across all generations produced by the model for the same prompt. In both methods, rewards are provided by a reward model. Specifically, for each input prompt, the model samples a set of $G$ candidate responses, forming a response group. The reward model assigns a score to each response, and advantages are computed by standardizing rewards within the group: subtracting the group mean and dividing by the group standard deviation. This yields relative, normalized advantage values. As in PPO, GRPO may incorporate a penalty term to prevent the updated policy from drifting too far from the reference policy, typically enforced via a Kullback–Leibler (KL) divergence regularizer to ensure stable updates. A formal definition of the GRPO objective, along with additional details, is provided in App. C.
4 Proposed Method: Murphy
Figure 2: Overview of Murphy. Given an input prompt ($q_{(\cdot)}$), $G_{1}$ code generations ($o_{(\cdot)}$) are generated and evaluated using a reward function ($r_{(\cdot)}$). Generations that do not achieve the maximum reward are revised based on executor feedback (f), combining the original prompt with the failed output, and re-prompted to generate another $G_{2}$ candidates. This iterative process continues for a fixed number of turns, with rewards from later turns propagated backward. The example illustrates the case with $G_{1}=G_{2}=2$, where $G_{1},G_{2}$ represents the number of rollouts per prompt, per turn, and $\rho(\cdot)$ denotes the credit assignment strategy.
Extending GRPO to multi-turn interaction requires formalizing how environmental feedback informs optimization. We study a feedback-rich code generation setting where model outputs are executed and scored against test suites, yielding two feedback types: (1) quantitative (e.g., proportion of test cases passed) and (2) qualitative (e.g., error traces or failing cases). Murphy extends GRPO by introducing (i) a multi-turn rollout mechanism that conditions generation on feedback, and (ii) a credit-assignment scheme that propagates rewards from future successful turns to earlier attempts.
Multi-turn rollout: For each task prompt, the policy begins at turn 1 by generating $G_{1}$ candidate solutions, each evaluated for reward against the prompt’s test suite. Failed generations are paired with their corresponding feedback and, together with the original prompt and prior outputs, form new conditioning contexts for subsequent turns. At turn $s$, the model generates $G_{s}$ new candidates for each failed case, continuing this feedback–refinement process for a fixed number of turns. This per-prompt, iterative rollout enables progressive improvement conditioned on feedback. Unlike standard GRPO, Murphy delays advantage computation until rollouts in all turns are generated and processed, allowing final rewards to retroactively shape credit assignment across prior turns.
Credit assignment: Once rewards for all turns are computed, Murphy backpropagates rewards from the final turn to earlier turns using a temporal credit-assignment scheme. Although early generations may achieve low initial rewards, incorporating execution feedback in later turns often leads to successful refinements. To ensure that these intermediate steps are properly credited, rewards from later successful generations are distributed to earlier ones according to their contribution to eventual success. After reward redistribution, advantages are computed for each generation, and the Murphy loss is applied to update the policy. Together, these mechanisms extend GRPO to handle multi-turn, feedback-conditioned optimization. Below, we formalize the multi-turn rollout mechanism and credit-assignment process and define the Murphy objective, extending GRPO to incorporate multi-turn feedback.
Notation and formalism.
We denote the current and old policy models by $\pi_{\theta}(\cdot\mid\cdot)$ and $\pi_{\theta_{\text{old}}}(\cdot\mid\cdot)$. $\pi_{\text{ref}}$ denotes the reference model (i.e., the model prior to fine-tuning), which is kept fixed (not updated) throughout training. $\mathcal{P}(Q)$ denotes the distribution over input prompts/questions $Q$, and $O$ the output space. As described earlier, in the first turn, the model generates $G_{1}$ candidate solutions for each prompt. For generations that do not attain the max reward, feedback is obtained from the environment. The model is then re-prompted with the original prompt, its previous output, and the corresponding feedback to produce $G_{s}$ new generations per prompt at turn $s$. This iterative procedure naturally forms a tree structure, where the root corresponds to the original prompt, and each subsequent generation (augmented with feedback) forms a child node, with generations at turn $s+1$ linked to their parent at turn $s$ (see Fig. 2).
Multi-turn rollout formalism.We define a feedback-conditioned rollout tree that captures how generations evolve over $S$ turns. Let $s$ denote the turn index and $G_{s}$ the number of generations per prompt at turn $s$. We use $i_{[1:s]}=(i_{1},\dots,i_{s})$ to index a path in the rollout tree, where $i_{j}$ denotes the branch taken at turn $j$.
Turn 1:The model receives $q_{(1)}\sim\mathcal{P}(Q)$ and generates a response group $\{o_{\{q_{(1)},j\}}\}_{j=1}^{G_{1}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{(1)})$. Each generation is executed to obtain a reward and feedback $(r_{\{q_{(1)},j\}},f_{\{q_{(1)},j\}})$, where $r$ denotes the proportion of tests passed and $f$ contains qualitative executor feedback (e.g., failing tests, error messages).
Turn s+1s{+}1:For any unsolved node at turn $s$ (i.e., one that does not achieve the maximum reward), we form a feedback-conditioned prompt by concatenation: $q_{(s+1,i_{[1:s]})}=[\,q_{(s,i_{[1:s-1]})},\,o_{\{q_{(s,i_{[1:s-1]})},i_{s}\}},\,f_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}\,],$ and sample $G_{s+1}$ children
$\{o_{\{q_{(s+1,i_{[1:s]})},k\}}\}_{k=1}^{G_{s+1}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{(s+1,i_{[1:s]})}),$
each evaluated to obtain $(r_{\{q_{(s+1,i_{[1:s]})},k\}},f_{\{q_{(s+1,i_{[1:s]})},k\}})$. This defines a rollout tree rooted at $q_{(1)}$ (see Fig. 2).
A complete path from the root to a leaf at turn $S$ can be written as
$q_{(1)}\rightarrow o_{\{q_{(1)},i_{1}\}}\rightarrow q_{(2,i_{[1:1]})}\rightarrow\cdots\rightarrow o_{\{q_{(S,i_{[1:S-1]})},i_{S}\}}.$
Leaf nodes at turn $S$ represent the final generations after refinement. We provide a more detailed formal treatment (including explicit Turn 2 expansion and tree construction details) in App. B.
Credit assignment formalism.After all outputs and corresponding rewards are generated and the rollout tree is constructed, we focus on assigning credit from later turns back to earlier ones. To achieve this, we explore two distinct strategies.
Max Reward Strategy (MaRS):The Max Reward Strategy is defined recursively, proceeding from the final turn back to the root. Since the final turn $S$ has no children, the rewards at this turn remain unchanged. We then consider turn s=S1s=S-1. Let $o_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}$ denote a generation at turn s=S1s=S-1 with an associated reward $r_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}$. If this generation already achieves the maximum reward, it has no children; otherwise, its children are defined as:
$C\big(o_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}\big)=\{o_{\{q_{(s+1,i_{[1:s]})},1\}},\ \ldots,\$
$o_{\{q_{(s+1,i_{[1:s]})},G_{S}\}}\}$
The corresponding set of rewards are defined as
$C_{r}\big(o_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}\big)=\{r_{\{q_{(s+1,i_{[1:s]})},1\}},\ \ldots,\$
$r_{\{q_{(s+1,i_{[1:s]})},G_{S}\}}\}$
which defaults to zero if there are no children. We then update the reward as:
$r_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}$ $=\text{max}\Bigg(r_{\{q_{(s,i_{[1:s-1]})},i_{s}\}},\$
$\quad\text{max}\Big(C_{r}(o_{\{q_{(s,i_{[1:s-1]})},i_{s}})\Big)\Bigg)$
This strategy assigns each node the maximum of its own reward and the best reward among its descendants. Intuitively, the descendant maximum represents the best outcome achievable through refinement, while taking the outer maximum ensures a node’s credit never decreases; even when feedback fails to improve performance. This formulation captures the maximum progress achievable from any refinement path starting at that node. The procedure operates recursively in a backward pass: rewards are first updated for all nodes at turn S1S-1 based on their children’s values at turn $S$. This process continues backward through the tree, with each turn $s$ receiving updated rewards based on the values from turn $s+1$, until reaching the root.
Mean Reward Strategy (MeRS):The Mean Reward Strategy follows the same recursive credit assignment structure as the Max Reward Strategy (MaRS), but differs in how rewards are propagated. Inspired by the return computation in REINFORCE (Williams, 1992), MeRS updates each node’s reward by incorporating the discounted mean of its children’s rewards.
For a generation $o_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}$ at turn $s$, let $\mathbb{I}_{\text{unsolved}}$ be an indicator that equals 1 if the problem remains unsolved at this node, and 0 otherwise. The update rule is:
$r_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}$ $=$
$\frac{r_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}+\gamma\cdot\overline{C}_{r}\Big(o_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}\Big)}{\mathbb{I}_{\text{unsolved}}\cdot(S-s)+1}$
where $\gamma\in[0,1]$ is a discount factor controlling the influence of descendant rewards, and $\overline{C}_{r}(\cdot)$ denotes the mean reward over children of unsolved nodes (children of solved nodes are masked out since they represent terminal states).
The denominator equals Ss+1S-s+1 for unsolved nodes and $1$ for solved nodes. This depth-based normalization serves two purposes: (1) it prevents rewards from growing unboundedly during backward propagation, as unsolved nodes can potentially accumulate discounted contributions from up to SsS-s future turns; and (2) it ensures fair comparison between nodes that solve at different turns, a node that solves immediately at turn $s$ retains its full reward, while a node that fails but has successful descendants receives appropriately scaled credit that accounts for the additional refinement steps required. All other aspects of the recursive procedure remain identical to MaRS.
MaRS vs MeRS:MaRS propagates the maximum descendant reward to the earlier turns, emphasizing peak performance, whereas MeRS propagates the mean reward, emphasizing stability and overall consistency. MaRS captures best-case improvement, while MeRS provides a smoother estimate of expected progress. Together, they offer complementary views of feedback-driven credit assignment.
Murphy Objective:Once the rewards are reassigned according to the chosen credit assignment strategy (MaRS or MeRS), advantages are computed similar to standard GRPO. The adjusted rewards serve as the basis for computing normalized advantages at each turn. Conditioned on a prompt $\tilde{q}$ and for $G_{s}$ generations, we normalize each reward by subtracting the mean reward and dividing by the standard deviation of the rewards obtained across the $G_{s}$ generations, yielding the normalized advantage $\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t}$. Additionally, as defined earlier, each $i$-th generation at turn $s$, $o_{\{q_{(s,i_{[1:s-1]})},i\}}\in O$, corresponds to a complete output trajectory, where $o_{\{q_{(s,i_{[1:s-1]})},i,t\}}$ denotes the $t$-th token and $o_{\{q_{(s,i_{[1:s-1]})},i,<t\}}$ the prefix up to (but excluding) token $t$. We denote the sequence length of the $i$-th generation by $|o_{\{q_{(s,i_{[1:s-1]})},i\}}|$. At each turn, the GRPO objective is applied using these credit-adjusted advantages. This per-turn optimization allows Murphy to incorporate feedback from later turns into earlier updates, effectively extending GRPO to a multi-turn setting. Finally, the divergence between the current and reference policies is captured by $D_{\mathrm{KL}}(\pi_{\theta}\,\|\,\pi_{{\text{ref}}})$, computed over all tokens in the generated sequences. The resulting optimization objective, which integrates credit-assigned rewards, normalized advantages, and KL regularization at each turn, defines the Murphy objective, distinguishing it from the standard GRPO formulation. The full Murphy objective is presented in Sec. 4.
Definition 1. (Murphy Objective) $\mathcal{J}_{\textsc{Murphy}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{P}(Q)}[\sum_{s,i_{1},\ldots i_{S}}\mathcal{J}_{q_{(s,i_{[1:s]})}}(\theta)]$ Where the per prompt objective at each turn $s$ is: $\mathcal{J}_{q_{(s,i_{[1:s-1]})}}(\theta)$ $=\mathbb{E}_{\{o_{(\tilde{q},i)}\}_{i=1}^{G_{s}}\sim\pi_{\theta_{\text{old}}}(O|\tilde{q})}\Bigg$[ $\sum_{i=1}^{G_{s}}\frac{1}{G_{s}|o_{\{\tilde{q},i\}}|}\sum_{t=1}^{|o_{\{\tilde{q},i\}}|}\Bigg($ $\min\Big(R_{\theta}(\tilde{q},i,t)\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t},$ $\text{clip}\left(R_{\theta}(\tilde{q},i,t),\,1-\epsilon,\,1+\epsilon\right)\cdot$ $\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t}\Big)-\beta D_{\mathrm{KL}}(\pi_{\theta}\,\|\,\pi_{{\text{ref}}})\Bigg)\Bigg]$ $\text{with,~}\tilde{q}$ $=q_{(s,i_{[1:s-1]})}$ $R_{\theta}(\tilde{q},i,t)$ $=\frac{\pi_{\theta}(o_{\{\tilde{q},i,t\}}\mid\tilde{q},o_{\{\tilde{q},i,<t\}})}{\pi_{\theta_{\text{old}}}(o_{\{\tilde{q},i,t\}}\mid\tilde{q},o_{\{\tilde{q},i,<t\}})}$ And, $\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t}$ denotes the advantage.
Note.
The design of Murphy is broadly applicable across a range of RLVR algorithms, including PPO (Schulman et al., 2017) and various extensions of GRPO (Ahmadian et al., 2024; Yu et al., 2025; Yue et al., 2025). In this work, we focus on GRPO due to its strong empirical performance in aligning LLMs (DeepSeek-AI et al., 2025; Yang et al., 2025). Extending Murphy to other RLVR variants is conceptually straightforward, as it builds on the same underlying principles. While our experiments primarily focus on code generation, where rich, verifiable feedback is readily available, the framework can naturally extend to other domains such as mathematics or logical reasoning, provided suitable forms of feedback are accessible.
4.1 Pruning Strategies in Murphy
By default, Murphy sets $G_{s}=G$ for all turns, maintaining a fixed number of generations per prompt throughout the multi-turn process. However, this multi-turn setup introduces significant computational cost. In the worst case, when success is achieved only at the final turn $S$, the number of generations per prompt can grow exponentially to $G^{S}$, resulting in a large search tree and substantial memory overhead. This makes Murphycomputationally expensive to optimize. While system-level optimizations such as vLLM with paged attention and KV caching (Kwon et al., 2023) make the generation process relatively efficient, the optimization phase remains costly. Large-scale rollouts can quickly exhaust GPU memory, and batching schemes that treat each generated sample as an independent training batch tend to be prohibitively slow. To address these challenges, we introduce two pruning strategies that reduce the number of rollouts at each turn, thereby making Murphy computationally tractable without compromising performance. We describe these strategies in detail below.
Intra-Group Pruning (IntraP):In IntraP, pruning operates recursively: starting from the children of level S1S-1, each pruning step is followed by reward propagation as described in Sec. 4, and the process continues backward until reaching the root. This ordering, pruning before reward reassignment, ensures that only informative trajectories are retained for credit propagation across turns. At each turn $s$, given a pruning budget $b$, we retain only the $b$ trajectories conditioned on the prompt whose rewards contribute most to the total reward variance within that group. The remaining trajectories, along with all their descendants, are discarded, and the process proceeds recursively until the root is reached. This approach is inspired by Xu et al. (2025a); they demonstrate that retaining trajectories with the highest reward variance within a group can reduce optimization cost while maintaining performance comparable to GRPO. We extend this principle to the multi-turn setting of Murphy.
Inter-Group Pruning (InterP):In InterP, the goal is to prune entire groups of children corresponding to generations conditioned on a given prompt. Specifically, each generation at a turn has a group of children, and we decide how many of these groups to retain based on a pruning budget $b$. To rank the groups, we assign a score inspired by the UCB sampling strategy (Auer et al., 2002), defined as $\alpha_{1}\mu+\alpha_{2}\sigma$, where $\mu$ and $\sigma$ denote the mean and standard deviation of the rewards within each group of children. The groups are ranked in descending order of this score, and only the top $b$ groups are retained, with $b$ determining the available computational budget. The remaining groups and their descendants are discarded. Similar to IntraP, this process is applied recursively and proceeds backward from the latter turns to the earlier ones. At each level, pruning is performed first, followed by credit assignment. In our experiments, we set $\alpha_{1}=0$ and $\alpha_{2}=1$, which effectively prioritizes groups exhibiting higher reward variance. Intuitively, a higher standard deviation not only captures greater variability within a group but also indicates that the model is uncertain about that group’s performance. Such groups likely contain some generations with near-maximum rewards, making them valuable targets for further optimization. In contrast, groups where all rewards are uniformly high (near $1$) or uniformly low (near 0) provide limited learning signal, as the model either has already mastered or completely failed the underlying behavior. High-variance groups, therefore, represent the most informative regions for continued improvement.
Table 1: Performance of Qwen3-1.7B, Qwen3-4B, and OLMo-2-1124-7B-Instruct variants across evaluation benchmarks. The Rollouts column indicates the total number of generations across two turns. Murphy (Ours) is highlighted. $\Delta_{3}$ denotes the difference between the Iter-3 performance of the GRPO/Murphy-trained models and that of Base (Iter-3), within each model block; green indicates improvement (darker = larger gain), red indicates regression.
Model Rollouts HumanEval (%) MBPP (%) BigCodeBench (%)
Iter-1 Iter-3 $\Delta_{3}$ Iter-1 Iter-3 $\Delta_{3}$ Iter-1 Iter-3 $\Delta_{3}$
Qwen3-1.7B
[…]
Qwen3-4B
[…]
OLMo-2-1124-7B-Instruct
[…]
5 Experiments
In this section, we first provide an overview of the models and datasets used in our experiments, followed by a detailed description of the setup and results, including ablation studies. We provide implementation details in App. G.
Training Dataset: We fine-tune all models on 1,000 samples randomly drawn from the KodCode dataset (Xu et al., 2025b). This dataset was chosen due to its minimal overlap with evaluation benchmarks [^1] Refer to Section 3.2 of KodCode (Xu et al., 2025b) for details on contamination analysis..
Evaluation Datasets. We evaluate trained models on a suite of programming benchmarks including HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), BigCodeBench-Hard (Zhuo et al., 2025) and Aider Polyglot [^2] https://epoch.ai/benchmarks/aider-polyglot. For BigCodeBench-Hard, which does not include visible unit tests, we randomly sample two test cases from the full test suite to construct visible tests and keep the test suite intact.
Metrics and Evaluation Protocol. Reflexion (Shinn et al., 2023) is a widely adopted inference-time iterative framework that enhances reasoning by refining incorrect outputs through iterative feedback and self-generated reflections (see App. E for details). To assess reasoning refinement and self-correction, we integrate all models into the Reflexion framework and report pass@1 under two settings: (i) Single iteration, equivalent to standard input-output prompting, and (ii) Three iterations, where feedback from evaluating visible test cases is incorporated into subsequent prompts. The agent terminates once all visible tests pass or when the maximum iteration limit is reached. Final solutions are evaluated on hidden test cases, and the resulting pass@1 is reported. Each experiment is repeated three times, and we report the mean and standard deviation of pass@1 across runs.
5.1 Murphy Experiments
We evaluate three models, Qwen3 (1.7B, 4B) (Yang et al., 2025) and OLMo-2-1124-7B-Instruct (OLMo et al., 2025), each fine-tuned on 1,000 samples randomly drawn from the KodCode dataset (Xu et al., 2025b). Evaluation is conducted on the benchmark datasets described in Sec. 5 using the Reflexion framework. We report pass@1 performance under both single and multi-iteration settings, as summarized in Tab. 1. For all our tables, best results are bold and second best are underlined. Results are reported as pass@1 accuracy (% mean $\pm$ stdev).
Reflexion: Single-iteration setting. (Iter-1 in Tab. 1) Models trained with the GRPO objective consistently outperform their base counterparts, achieving notable gains; for instance, a $\sim 3\%$ improvement on HumanEval for OLMo-2-1124-7B-Instruct. Murphy achieves competitive or superior performance than GRPO in this setting.
Reflexion: Multi-iteration setting. (Iter-3 in Tab. 1) We repeat the experiments with three iterations in the Reflexion framework to assess self-correction and reasoning-refinement capabilities. Increasing the number of iterations leads to consistent performance improvements across all models and benchmarks. Models trained with Murphy surpass both GRPO-trained and base models, achieving absolute gains of up to $8\%$ over GRPO. These results demonstrate the effectiveness of multi-turn reflective optimization in enhancing reasoning refinement and self-correction.
Table 2: Ablation comparing Max Reward (MaRS) vs Mean Reward (MeRS) propagation. MaRS (Ours) is highlighted. $\Delta_{3}$ compares Iter-3 to MeRS ($\gamma=1$) Iter-3 within each model block (green indicates improvement; neutral indicates no gain).
Model & Strategy Rollouts HumanEval (%) MBPP (%) BigCodeBench (%)
Iter-1 Iter-3 $\Delta_{3}$ Iter-1 Iter-3 $\Delta_{3}$ Iter-1 Iter-3 $\Delta_{3}$
Qwen3-1.7B
[…]
OLMo-2-1124-7B-Instruct
[…]
Table 3: Comparison of Murphy and its pruned variants on evaluation benchmarks using Qwen3-1.7B. All variants generate 72 rollouts per query. The Updates column denotes the total number of gradient steps per query. $\Delta_{3}$ compares Iter-3 against the unpruned Murphy (MaRS) Iter-3; green indicates improvement (darker = larger gain), red indicates regression.
Model & Strategy Updates HumanEval (%) MBPP (%) BigCodeBench (%)
Iter-1 Iter-3 $\Delta_{3}$ Iter-1 Iter-3 $\Delta_{3}$ Iter-1 Iter-3 $\Delta_{3}$
[…]
Comparisons to execution-feedback RL methods.
[…]
Generalization to Multi-Turn Code Editing.
We evaluate generalization on the Aider Polyglot Benchmark [^3] https://github.com/Aider-AI/aider/tree/main/benchmark, which tests self-correction through execution feedback: models receive two attempts per problem, with unit test results provided as feedback after a failed first attempt. This setup closely mirrors Murphy’s training objective. Without additional training, we evaluate the same Qwen3-4B checkpoints from our main experiments (Tab. 1) following the official evaluation setup. The Murphy-trained model achieves 10.9% pass@1 and 21.6% pass@2, outperforming both the base (7.6% / 18.7%) and GRPO-trained (8.0% / 17.8%) models, confirming that Murphy’s gains transfer to external feedback-driven tasks.
Additional Experiments. Further analysis is provided in the appendices: effect of training data size (App.G.4.2), computational cost (App.F), comparison with naive multi-turn GRPO (App.G.4.1), and training beyond two turns (App.G.4.3). App. D confirms that Murphy’s gains generalize across different iterative scaffolds.
5.2 Ablation 1: MaRS vs. MeRS
As described in Sec. 4, we compare two reward propagation strategies: MaRS and MeRS. We train Qwen3-1.7B and OLMo-2-1124-7B-Instruct on 1,000 KodCode samples under each strategy and report results in Tab. 2. Across both models and multiple Reflexion iterations, MaRS consistently matches or surpasses MeRS, independent of the discount factor $\gamma$. The key difference lies in handling non-binary rewards: MeRS averages rewards across generations, diluting the learning signal when few outputs score highly, whereas MaRS propagates the strongest outcome, allowing rare but valuable high-reward trajectories to dominate the update. This makes MaRS particularly effective in multi-turn settings with sparse rewards, while the gap between the two strategies diminishes in binary-reward tasks.
5.3 Ablation 2: IntraP vs. InterP
In Subsec. 4.1, we introduced two pruning strategies: Intra-Group (IntraP) and Inter-Group (InterP) pruning. Both reduce the rollout tree size, thereby decreasing the total number of gradient updates (Tab. 3). We compare their effectiveness using Qwen3-1.7B. While both strategies reduce computational cost, they exhibit different trade-offs. IntraP improves single-turn performance (Iter-1) across all benchmarks but shows modest regression on multi-turn evaluation (Iter-3) for HumanEval and MBPP. In contrast, InterP maintains comparable multi-turn performance on HumanEval and MBPP while achieving notable gains on BigCodeBench. These results suggest that InterP offers a more robust trade-off, preserving multi-turn self-correction capabilities while reducing computational cost.
6 Conclusion & Limitations
We introduced Murphy, a multi-turn reflective reinforcement learning framework that extends RLVR algorithms by incorporating iterative self-correction through both quantitative and qualitative feedback. By grounding optimization in intermediate feedback signals and propagating rewards across refinement turns, Murphy consistently improves reasoning and code generation performance, particularly in multi-iteration settings where feedback-driven refinement is crucial. These findings underscore the value of integrating structured feedback directly into the optimization process. While effective, Murphy’s multi-turn design increases computational cost; pruning partially mitigates this but it remains more resource-intensive than single-turn baselines. We study structured execution feedback for code, and generalization to noisier feedback, deeper refinement, and broader agentic objectives is open. In future work, we aim to develop adaptive turn/rollout selection and tool-augmented optimization (retrieval, execution, APIs).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix
This appendix complements the main text with extended related work, formal definitions, and additional analyses and ablations. The contents are organized as follows:
  • App. A: Extended discussion of related work, covering LLM agents for software development and reinforcement learning with verifiable rewards.
  • App. B: Detailed treatment of the multi-turn rollout tree formalism and indexing conventions.
  • App. C: Review of the GRPO objective and associated notation.
  • App. D: Sensitivity analysis of Murphy-trained models to different multi-iteration inference scaffolds.
  • App. E: Description of the Reflexion framework used for evaluation.
  • App. F: Computational cost comparison between multi-turn and single-turn training.
  • App. G: Implementation details, hyperparameters, and additional experiments including ablations on dataset size, naive multi-turn baselines, and pruning strategies.
  • App. H: Prompts used in our experiments
  • App. I: Potential risks
  • App. J: Icon attributions for figures in the main text.
Appendix A Extended Related Work
LLM Agents for Software Development.
Recent works (Jiang et al., 2025; Zhong et al., 2024) have explored LLM agents for programming tasks such as code generation, bug fixing, and code migration. A key driver of progress in these domains has been inference-time iterative frameworks (Shinn et al., 2023; Lingam et al., 2025), which leverage execution feedback to generate self-reflections for refining candidate programs (Yang et al., 2024; Xia et al., 2025). While these approaches underscore the value of iterative feedback and scaffolding, they primarily enhance the inference pipeline rather than the underlying model. Our work takes a complementary direction: we improve the reasoning and self-correction abilities of LLMs themselves through training-time optimization, thereby strengthening the base models that agentic frameworks depend on.
Reinforcement Learning with Verifiable Rewards for LLM Reasoning.
[…]
Appendix B Multi-turn Rollout Formalism (Detailed)
This appendix section provides a detailed treatment of the rollout-tree construction and indexing used throughout the main paper.
We define a feedback-conditioned rollout tree that captures how model generations evolve across multiple turns of interaction with the environment. Let $s$ denote the turn index, $S$ the total number of turns, and $G_{s}$ the number of generations per prompt at turn $s$.
Turn 1:In the first turn ($s=1$), the model receives an input prompt $q_{(1)}$ sampled from $\mathcal{P}(Q)$ and generates $G_{1}$ candidate programs which forms a response group $\{o_{\{q_{(1)},1\}},o_{\{q_{(1)},2\}},\ldots,o_{\{q_{(1)},G_{1}\}}\}$. Note that $\{o_{\{q_{(1)},j\}}\}_{j=1}^{G_{1}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{(1)})$. Each generation is executed against its associated test suite, producing a numerical reward $r_{\{q_{(1)},j\}}$ and environment feedback $f_{\{q_{(1)},j\}}$. The reward represents the proportion of test cases passed. The environment feedback contains the specific unit tests that passed or failed, along with any corresponding error messages. These $G_{1}$ generations form the first layer of output nodes in the rollout tree.
Turn 2:For each generation $j$ in turn 1 that fails to achieve the maximum reward (which equals 1 in our setting, since it represents the proportion of test cases passed), the corresponding feedback is appended to the original prompt and the prior output from turn $1$ to form a feedback-conditioned prompt: $q_{(2,j)}=[\ q_{(1)},\ o_{\{q_{(1)},j\}},\ f_{\{q_{(1)},j\}}\ ]$ where $\cdot\,$] denotes textual concatenation. The model is then re-invoked to generate $G_{2}$ new candidate solutions: $\{o_{\{q_{(2,j)},k\}}\}_{k=1}^{G_{2}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{(2,j)})$. Each of these generations is evaluated to obtain $(r_{\{q_{(2,j)},k\}},f_{\{q_{(2,j)},k\}})$, which denote the reward and feedback. These output generations represent refinements of their corresponding parent outputs $o_{\{q_{(1)},j\}}$ and collectively form the second layer of the rollout tree.
Turn $s$:Building on the previous turn, this procedure extends recursively to any turn $s\in\{1,\dots,S-1\}$. We define $i_{[1:s]}=i_{1},\dots,i_{s}$ as a sequence of branch indices that trace a specific path through the tree, where $i_{j}$ indicates the $i$’th candidate that was selected at turn $j$. Similarly, $i_{[1:1]}$ denotes $i_{1}$. For each generation $o_{{q_{(s,i_{[1:s-1]})},i_{s}}}$ that fails to achieve the maximum reward, we construct a feedback-conditioned prompt:
$q_{(s+1,i_{[1:s]})}=[q_{(s,i_{[1:s-1]})},\,o_{{q_{(s,i_{[1:s-1]})},i_{s}}},\,f_{{q_{(s,i_{[1:s-1]})},i_{s}}}]$
where $\cdot$] denotes textual concatenation. The model is then re-invoked to generate $G_{s+1}$ new candidate solutions:
$\{o_{\{q_{(s+1,i_{[1:s]})},k\}}\}_{k=1}^{G_{s+1}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{(s+1,i_{[1:s]})})$
Each candidate is evaluated to obtain its reward and feedback $(r_{\{q_{(s+1,i_{[1:s]})},k\}},\,f_{\{q_{(s+1,i_{[1:s]})},k\}})$. The resulting generations, $\{o_{\{q_{(s+1,i_{[1:s]})},k\}}\}_{k=1}^{G_{s+1}}$, form the child nodes of parent node $o_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}$.
A complete path from the root (the initial prompt at turn 1) to a leaf at turn $S$ (final output) can be expressed as:
$q_{(1)}\rightarrow o_{{q_{(1)},i_{1}}}\rightarrow q_{(2,i_{[1:1]})}\rightarrow o_{{q_{(2,i_{[1:1]})},i_{2}}}\rightarrow\cdots\rightarrow o_{\{q_{(S,i_{[1:S-1]})},i_{S}\}}$
where the indices $i_{1},\dots,i_{S}$ specify which branch is taken at each turn. Leaf nodes at turn $S$ represent the final generations obtained after completing all refinement steps. Once rewards for all turns are computed, the rewards from these terminal nodes are propagated backward through their ancestors according to the credit assignment strategies described below.
Appendix C GRPO: Objective and Additional Details
Notation.
We denote the model policy by $\pi_{\theta}(\cdot\mid\cdot)$ and the reference (older) policy by $\pi_{\theta_{\text{old}}}(\cdot\mid\cdot)$. Let $G$ be the number of generations per prompt, $\mathcal{P}(Q)$ the distribution over input prompts/questions $Q$, and $O$ the output space. For a given prompt $q\sim\mathcal{P}(Q)$, the reference policy produces a set of $G$ responses, forming a response group $\{o_{q,1},o_{q,2},\ldots,o_{q,G}\}$. Each generation $o_{q,i}\in O$ corresponds to a full output trajectory, where $o_{q,i,t}$ denotes the $t$-th token and $o_{q,i,<t}$ the prefix up to (but excluding) token $t$. We write $|o_{q,i}|$ for the sequence length of the $i$-th generation. For a given prompt $q$, the reward model assigns a scalar score to each response in the group, yielding $\mathbf{r}_{q}=\{r_{q,1},r_{q,2},\ldots,r_{q,G}\}$. Moreover, for prompt $q$, the advantage associated with the $t$-th token of the $i$-th generation is defined as $\hat{A}_{q,i,t}=(r_{q,i}-\mu(\mathbf{r}_{q}))/\sigma(\mathbf{r}_{q})$, where $\mu(\mathbf{r}_{q})$ and $\sigma(\mathbf{r}_{q})$ denote the mean and standard deviation of the group rewards, respectively. Finally, $D_{\mathrm{KL}}(\pi_{\theta}||\pi_{\theta_{\text{old}}})$ denotes the KL divergence between the current and reference policies, computed over all tokens in the generated sequences. The GRPO training objective is presented in App. C.
Definition 2. (GRPO Objective) $\mathcal{J}(\theta)=\$$\mathbb{E}_{q\sim\mathcal{P}(Q),\ \{o_{(q,i)}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}\Big[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{q,i}|}\sum_{t=1}^{|o_{q,i}|}\min\Big(R_{\theta}(q,i,t)\hat{A}_{q,i,t},$ $\text{clip}\Big(R_{\theta}(q,i,t),\,1-\epsilon,\,1+\epsilon\Big)\hat{A}_{q,i,t}\Big)\Big]-\beta D_{\text{KL}}(\pi_{\theta}\,\|\,\pi_{\text{ref}})$ Where, $R_{\theta}(q,i,t)$ $=\frac{\pi_{\theta}(o_{q,i,t}\mid q,o_{q,i,<t})}{\pi_{\theta_{\text{old}}}(o_{q,i,t}\mid q,o_{q,i,<t})}$ $\hat{A}_{q,i,t}$ $=\frac{(r_{q,i}-\mu(\mathbf{r}_{q}))}{\sigma(\mathbf{r}_{q})}$
Appendix D Sensitivity to Multi-Iteration Scaffolds
Murphy is explicitly designed to train models to incorporate environment feedback and self-correct over multiple turns. In a single-iteration evaluation setting, where no feedback is available, Murphy-trained models naturally perform comparably to GRPO-trained models. In contrast, under multi-iteration evaluation—typical of agentic settings where feedback is present—Murphy yields substantially larger gains. This behavior is expected and reflects the objective of Murphy, which is to improve a model’s ability to utilize feedback across turns rather than to optimize single-pass generation.
To assess sensitivity to the choice of multi-iteration scaffold, we conduct additional experiments using the Qwen3-1.7B model. We compare the base model, the GRPO-trained model, and the Murphy-trained model under three distinct scaffolds: Reflexion (Shinn et al., 2023), LATS (Zhou et al., 2024), and DoT (Lingam et al., 2025). Importantly, no additional training is performed. We evaluate the same GRPO checkpoint reported in Table 1 and the same Murphy (MaRS, IntraP) checkpoint reported in Table 3. All scaffolds use their standard configurations with three iterations. We report mean accuracy and standard deviation over three independent runs on HumanEval.
Table 4: Sensitivity of Qwen3-1.7B performance on HumanEval to different multi-iteration scaffolds. No additional training is performed.
Model Reflexion LATS DoT
Qwen3-1.7B (Base) 80.07 $\pm$ 1.27 78.95 $\pm$ 1.29 78.35 $\pm$ 3.88
Qwen3-1.7B GRPO 82.11 $\pm$ 1.76 80.79 $\pm$ 3.01 83.74 $\pm$ 0.57
Qwen3-1.7B Murphy (MaRS) 84.28 $\pm$ 2.01 86.89 $\pm$ 0.42 85.66 $\pm$ 0.43
Across all scaffolds, the Murphy-trained model consistently outperforms both the GRPO-trained model (trained under a matched compute budget) and the base model. These results indicate that the gains from Murphy are not dependent on any particular inference-time scaffold, but instead reflect improved feedback utilization learned during training.
Appendix E Reflexion
Reflexion (Shinn et al., 2023) is an inference-time iterative framework designed to improve reasoning through repeated interaction with feedback from an external environment. It employs three agents: an actor ($M_{a}$), an evaluator ($M_{e}$), and a self-reflection module ($M_{sr}$), which operate cyclically until a termination condition is met. For code generation, the process proceeds as follows:
  1. 1.
    Actor step: The actor $M_{a}$ receives an input and generates an output (e.g., a code snippet).
  2. 2.
    Evaluation step: The evaluator $M_{e}$ scores the output (e.g., the percentage of unit tests passed).
  3. 3.
    Self-reflection step: If the score is insufficient, the self-reflection module $M_{sr}$ diagnoses the issue, proposes a fix, and appends both the failed output and the suggested correction to the input context. The updated input is then fed back to the actor, and the cycle repeats until either the task succeeds or a maximum number of iterations is reached.
In most implementations, the actor and the self-reflection module are instantiated by the same underlying language model. The self-reflection stage thus corresponds to the model reasoning over its own prior outputs augmented with feedback from the evaluator (or executor) and the previous input, output pairs, to generate improved responses in subsequent iterations.
Appendix F Computational Cost of Multi-Turn vs. Single-Turn GRPO
We compare the computational cost of multi-turn Murphy training against single-turn GRPO under matched rollout and gradient budgets (72). Table 5 reports end-to-end FLOPs per GPU and per-step training latency computed using DeepSpeed’s (Rasley et al., 2020) flops-profiler [^4] https://www.deepspeed.ai/tutorials/flops-profiler/ for Qwen3-1.7B and Qwen3-4B. These measurements capture the full optimization step, including forward and backward passes, optimizer updates, and communication overhead.
In some cases, multi-turn Murphy may require fewer FLOPs per step than single-turn GRPO despite incorporating feedback into later-turn prompts. This reduction arises from the tree structure of multi-turn rollouts: only generations that fail to solve the problem are expanded in subsequent turns, while successful generations terminate early. Under a matched rollout budget, this early termination reduces the total number of tokens generated, outweighing the increased sequence length from feedback concatenation.
However, multi-turn training incurs moderately higher step latency (e.g., 1.34s vs. 1.15s for Qwen3-1.7B). This is because later-turn sequences are longer due to feedback concatenation, leading to increased memory bandwidth requirements and quadratically scaling attention costs even though fewer sequences are processed overall. The per-device batch size is set to 8 for Qwen3-1.7B and 4 for Qwen3-4B.
Table 5: End-to-end computational cost per optimization step for multi-turn Murphy and single-turn GRPO under matched rollout and gradient budgets.
Model Setting FLOPs / GPU Step Latency
Qwen3-1.7B Multi-turn (Murphy) 23.39 TFLOPs 1.34 s
Single-turn (GRPO) 27.69 TFLOPs 1.15 s
Qwen3-4B Multi-turn (Murphy) 10.49 TFLOPs 2.34 s
Single-turn (GRPO) 11.68 TFLOPs 2.20 s
Appendix G Implementation Details
We implement our framework on top of TRL (von Werra et al., 2020), which provides efficient distributed training and a modular implementation of GRPO. We integrate TRL with vLLM for fast inference and large-scale rollout execution, enabling scalable multi-turn training in our experiments. Prompts used to train Murphy are listed in App. H. All experiments use publicly available datasets. The base models (Qwen3, OLMo) are available for research use under the Apache 2.0 license.
G.1 Model Size and Compute Budget
All experiments were conducted on $8\times$ NVIDIA H100 GPUs. Our implementation builds on HuggingFace’s TRL [^5] https://huggingface.co/docs/trl/en/index. For efficiency, 2 GPUs were allocated for inference via vLLM, while the remaining 6 GPUs handled model updates. Training Qwen3-1.7B with Murphy on 1,000 KodCode samples took approximately 1.5 hours, Qwen3-4B took 4 hours, and OLMo-2-7B-Instruct required 10 - 13 hours. Checkpoints were saved every 50 steps, and for all baselines, we selected the checkpoint corresponding to one epoch.
G.2 Hyperparameters
We set the KL regularization factor $\beta=0.04$, learning rate to $10^{-6}$, and weight decay to $0.1$ for both GRPO and Murphy  variants. Unless stated otherwise, the number of turns in Murphy is set to 2. For GRPO and the first turn of Murphy, we use 8 rollouts per prompt, while the second turn uses up to 8 rollouts per prompt (resulting in a maximum of 64 rollouts in turn 2). To ensure a fair computational comparison, we also train GRPO with 72 rollouts. Following Yang et al., we set the temperature to 0.6 and top-$p$ to 0.95 during inference for all experiments.
G.3 Package Parameters
We use the Reflexion (Shinn et al., 2023) framework to evaluate all trained models. The number of iterations is swept over $\{1,3\}$, and max-tokens is set to each model’s maximum generation length. Models are hosted via vLLM. Since all models fit on a single H100 GPU, we set data-parallel-size to 8 and enable prefix caching to accelerate evaluation. We use the following commands to install the appropriate packages:
pip install uv && \
uv pip install trl==0.19.1 && \
uv pip install gunicorn==20.1.0 && \
uv pip install fastapi==0.115.12
uv pip install uvicorn==0.34.2 && \
uv pip install aiohttp==3.11.18
uv pip install astunparse==1.6.3
uv pip install jsonlines tenacity && \
uv pip install vllm==0.8.5.post1
G.4 Additional Experiments
Table 6: Performance of Qwen3-1.7B and OLMo-2-1124-7B-Instruct variants on HumanEval, MBPP, and BigcodeBench benchmarks, reported as pass@1 accuracy (% mean $\pm$ stdev over 3 independent runs). Best results are bold, second-best are underlinedMurphy- MaRS outperforms Murphy (Simple) on average. Murphy (Simple) is “simple” in that it uses no fan-out: it performs 144 total rollouts by sampling 72 outputs in the first turn and then sampling another 72 outputs in a second turn that incorporates feedback from the first turn. We match the total number of rollouts/updates to ensure a fair comparison.
Model Rollouts HumanEval (%) MBPP (%) BigCodeBench (%)
Iter-1 Iter-3 Iter-1 Iter-3 Iter-1 Iter-3
Qwen3-1.7B
     Base 74.19 $\pm$ 1.95 80.07 $\pm$ 1.27 43.47 $\pm$ 0.64 53.93 $\pm$ 3.52 7.20 $\pm$ 2.73 18.24 $\pm$ 1.79
     GRPO 72 77.42 $\pm$ 0.59 82.11 $\pm$ 1.76 45.93 $\pm$ 1.60 57.20 $\pm$ 1.40 10.58 $\pm$ 1.95 20.49 $\pm$ 2.17
     Murphy (Simple) 144 70.33 $\pm$ 0.93 82.11 $\pm$ 0.35 46.00 $\pm$ 0.40 58.53 $\pm$ 1.03 13.29 $\pm$ 1.41 18.24 $\pm$ 1.17
     Murphy- MaRS 72 79.67 $\pm$ 3.01 86.58 $\pm$ 1.06 44.73 $\pm$ 0.50 62.00 $\pm$ 1.91 6.98 $\pm$ 3.72 20.25 $\pm$ 3.07
OLMo-2-1124-7B-Instruct
     Base 37.20 $\pm$ 0.86 46.04 $\pm$ 0.43 19.90 $\pm$ 0.42 29.60 $\pm$ 0.28 1.35 $\pm$ 0.00 3.72 $\pm$ 0.48
     GRPO 72 41.26 $\pm$ 0.35 43.70 $\pm$ 0.93 32.87 $\pm$ 0.23 35.80 $\pm$ 1.00 0.68 $\pm$ 0.00 2.70 $\pm$ 0.68
     Murphy (Simple) 144 39.02 $\pm$ 0.00 41.87 $\pm$ 0.93 28.33 $\pm$ 0.31 33.60 $\pm$ 0.69 1.13 $\pm$ 1.41 2.25 $\pm$ 0.39
     Murphy- MaRS 72 45.53 $\pm$ 0.70 52.24 $\pm$ 1.96 29.33 $\pm$ 0.90 39.67 $\pm$ 1.29 1.80 $\pm$ 0.39 3.38 $\pm$ 1.17
G.4.1 Ablation: Comparison with a Naive Multi-Turn Extension
A straightforward way to adapt GRPO for multi-turn training is to extend failed rollouts by appending feedback from previous turns without introducing additional fan-out. Rewards and advantages are then computed on the final turn, and the GRPO objective is applied. We compare this baseline against Murphy in Tab. 6. The number of turns is set to 2. The results underscore the importance of Murphy’s structured credit assignment and staged fan-out, both of which contribute to its superior multi-turn optimization performance.
Table 7: Ablation: average $\pm$ stdev (percent) for Murphy vs GRPO on KodCode (2K/3K/4K). Murphy MaRS shows gains of up to $\sim 9\%$ over compute equivalent GRPO.
Qwen3-1.7B Dataset HumanEval HumanEval MBPP MBPP BigcodeBench BigcodeBench
Iter-1 Iter-3 Iter-1 Iter-3 Iter-1 Iter-3
Murphy- MaRS 2K $\mathbf{81.65\%\pm 2.19}$ $\mathbf{84.15\%\pm 2.05}$ $\mathbf{46.93\%\pm 0.81}$ $\mathbf{60.47\%\pm 2.66}$ $\mathbf{2.70\%\pm 0.68}$ $\mathbf{11.04\%\pm 1.70}$
GRPO 2K $78.46\%\pm 2.46$ $83.54\%\pm 1.06$ $46.13\%\pm 0.70$ $57.67\%\pm 0.61$ $2.48\%\pm 0.78$ $8.11\%\pm 1.17$
Murphy- MaRS 3K $77.30\%\pm 3.55$ $\mathbf{84.55\%\pm 1.96}$ $45.53\%\pm 0.12$ $\mathbf{63.60\%\pm 0.72}$ $\mathbf{3.83\%\pm 0.78}$ $\mathbf{10.36\%\pm 1.41}$
GRPO 3K $\mathbf{79.07\%\pm 2.54}$ $80.49\%\pm 1.61$ $\mathbf{47.40\%\pm 1.20}$ $54.87\%\pm 1.68$ $3.60\%\pm 1.03$ $6.08\%\pm 0.68$
Murphy- MaRS 4K $79.27\%\pm 2.26$ $\mathbf{89.63\%\pm 1.22}$ $45.60\%\pm 1.31$ $\mathbf{62.80\%\pm 0.92}$ $\mathbf{3.60\%\pm 0.39}$ $\mathbf{8.78\%\pm 2.70}$
GRPO 4K $\mathbf{79.88\%\pm 2.66}$ $83.33\%\pm 0.70$ $\mathbf{47.27\%\pm 0.46}$ $57.40\%\pm 1.83$ $2.70\%\pm 1.17$ $6.98\%\pm 1.41$
G.4.2 Ablation: Effect of Training Dataset Size
To examine the effect of training dataset size, we construct three nested subsets of KodCode ($2\subset$ $3\subset$ 44K) and train Qwen3-1.7B using GRPO (72 rollouts) and Murphy. Results are summarized in Tab. 7. While performance does not increase monotonically with dataset size, Murphy consistently demonstrates superior multi-turn robustness and maintains an average improvement of up to $\sim 9\%$ over GRPO across all dataset scales.
G.4.3 Ablation: Effect of Pruning in Multi-Turn Murphy Training
We noted in the main paper that increasing the number of turns in Murphy can lead to exponential growth in computational cost. To mitigate this, we design and evaluate two pruning strategies. In this experiment, we extend our setup to a 3-turn setting and study the effect of pruning on performance. Results in Tab. 8 show that the pruned variant achieves competitive or even superior performance compared to the non-pruned counterpart.
Table 8: Ablation study comparing pruning versus non-pruning strategies for Qwen3-1.7B trained with Murphy for 3 turns on the KodCode dataset. Reported numbers indicate pass@1 (%) over three independent runs. Pruned variant achieves competitive or superior performance compared to non-pruned Murphy.
Model HumanEval MBPP BigCodeBench
Iter-1 Iter-3 Iter-1 Iter-3 Iter-1 Iter-3
GRPO $77.24\pm 2.14$ $83.13\pm 1.96$ $44.07\pm 0.31$ $53.93\pm 0.61$ $1.13\pm 0.39$ $6.53\pm 0.78$
Murphy- MaRS, InterP $78.66\pm 2.66$ $\mathbf{84.15\pm 1.83}$ $\mathbf{45.53\pm 1.17}$ $61.20\pm 0.53$ $\mathbf{2.70\pm 0.68}$ $\mathbf{7.43\pm 1.79}$
Murphy- MaRS $\mathbf{79.27\pm 2.11}$ $83.94\pm 1.41$ $44.80\pm 1.11$ $\mathbf{62.00\pm 1.40}$ $1.35\pm 0.68$ $5.86\pm 0.39$
Appendix H Prompt Examples
To train the Murphy objective, we employ the following prompts. The system prompt is used at each dialogue turn, and the feedback prompts are applied during every feedback turn.
System Prompt You are a Python coding AI agent. When given a Python function signature and docstring, you must provide a complete Python solution following this exact format: 1. Reasoning Phase: Use <think>...</think> tags to contain your complete thought process. Break down the problem requirements step by step Identify key constraints, edge cases, and potential pitfalls Plan your algorithm and data structures Walk through examples to validate your approach Explain your logic thoroughly as if working on scratch paper 2. Implementation Phase: Use <output>...</output> tags to contain your final code. Include the complete function with the original signature Ensure your code directly implements the approach from your thinking Write clean, readable code with appropriate comments if needed Your solution must be complete, correct, and handle all specified requirements.
Feedback Header You have previously attempted this problem {num-attempts} time(s).
Feedback Body Previous Attempt #{num-attempts} Analysis: Your earlier thought process and implementation: Let me think step by step. <think> {code-generated} Test Case Results: - Passed: {passed-tests}/{total-tests} test cases - Failed: {failed-tests}/{total-tests} test cases Detailed Feedback: {feedback-string} Note: For failed test cases, your code’s output is shown followed by a ‘#‘ symbol indicating the failure.
Feedback Footer. Instructions for Your Next Attempt: 1. Failure Analysis Phase: (a) Carefully examine each failed test case to understand exactly what went wrong (b) Identify the specific lines of code or logic that caused the failures (c) Look for patterns across multiple failed cases (e.g., all involve negative numbers, empty inputs, etc.) (d) Determine if failures stem from algorithmic errors, edge case handling, or implementation bugs 2. Solution Refinement Phase: (a) Build upon any correct aspects of your previous solution that passed tests (b) Redesign the problematic parts of your algorithm to handle the failed cases (c) Ensure your new approach covers edge cases that weren’t properly handled before (d) Consider additional edge cases that might not be in the test suite but could break your solution 3. Implementation Phase: (a) Write your improved solution that directly addresses the identified failure points (b) Test your logic mentally against the failed cases to verify it would now pass (c) Ensure your solution maintains correctness for previously passing test cases Use the concrete feedback from your {num-attempts} previous attempt(s) to create a more robust and accurate solution.
Appendix I Potential Risks
While Murphy introduces some new dynamics through iterative self-correction and reflective optimization, the associated risks appear modest overall. The main considerations involve ensuring that feedback loops remain interpretable and that reward signals do not inadvertently reinforce narrow or heuristic reasoning. There is also some potential for subtle reward hacking, where the model optimizes for easily verifiable but shallow improvements, or for mild distributional drift if reflective heuristics fail to generalize beyond training contexts. Nonetheless, because Murphy still relies on verifiable rewards and bounded reflection, these risks are relatively contained and can be mitigated through careful evaluation design, human oversight, and robust validation across diverse task settings.
Appendix J Icon Attributions